### v4

# On semipositone problems over ${\mathbb{R}}^{N}$ for the fractional $p$-Laplace operator

Abstract.
For $N\geq 1,s\in(0,1)$, and $p\in(1,\frac{N}{s})$ we find a positive solution to the following class of semipositone problems associated with the fractional $p$-Laplace operator:
$(-\Delta)_{p}^{s}u=g(x)f_{a}(u)\text{ in }\mathbb{R}^{N},$ (SP)
where $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ is a positive function, $a>0$ is a parameter and $f_{a}\in{\mathcal{C}}(\mathbb{R})$ is defined as $f_{a}(t)=f(t)-a$ for $t\geq 0$, $f_{a}(t)=-a(t+1)$ for $t\in[-1,0]$, and $f_{a}(t)=0$ for $t\leq-1$, where $f$ is a non-negative continuous function on $[0,\infty)$ satisfies $f(0)=0$ with subcritical and Ambrosetti-Rabinowitz type growth. Depending on the range of $a$, we obtain the existence of a mountain pass solution to (SP) in $\mathcal{D}^{s,p}(\mathbb{R}^{N})$. Then, we prove mountain pass solutions are uniformly bounded with respect to $a$, over $L^{r}(\mathbb{R}^{N})$ for every $r\in\left[\frac{Np}{N-sp},\infty\right]$. In addition, if $p>\frac{2N}{N+2s}$, we establish that (SP) admits a non-negative mountain pass solution for each $a$ near zero. Finally, under the assumption $g(x)\leq\frac{B}{|x|^{\beta(p-1)+sp}}$ for $B>0,x\neq 0$, and $\beta\in\left(\frac{N-sp}{p-1},\frac{N}{p-1}\right)$, we derive an explicit positive radial subsolution to (SP) and show that the non-negative solution is positive a.e. in $\mathbb{R}^{N}$.
Key words and phrases:
semipositone problems; fractional $p$-Laplace operator; uniform regularity estimates; positive solutions.
2020 Mathematics Subject Classification:
35D30, 35A15, 35R11, 35B65, 35B09
2Corresponding author
1. Introduction
In this article, for $N\geq 1,s\in(0,1)$, and $p\in(1,\frac{N}{s})$ we study the following semipositone problems associated with the fractional $p$-Laplace operator:
$(-\Delta)_{p}^{s}u=g(x)f_{a}(u)\text{ in }{\mathbb{R}}^{N},$ (SP)
where the function $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ is positive, $a>0$ is a parameter and the associated function $f_{a}:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is defined as follows:
$f_{a}(t)=\begin{cases}f(t)-a&\text{ if }t\geq 0,\\ -a(t+1)&\text{ if }t\in[-1,0],\\ 0&\text{ if }t\leq-1,\end{cases}$ (1.1)
where $f$ is a non-negative continuous function on $[0,\infty)$ with $f(0)=0$. Also, $f$ satisfies the following growth assumptions:
  1. ($\bf f{1}$)
    $\lim\limits_{t\rightarrow 0^{+}}\frac{f(t)}{t^{p-1}}=0,$ and $\lim\limits_{t\rightarrow\infty}\frac{f(t)}{t^{\gamma-1}}\leq C(f)$ for some $\gamma\in\left(p,\frac{Np}{N-sp}\right)$ and $C(f)>0$,
  2. ($\bf f{2}$)
    (Ambrosetti-Rabinowitz) there exist $\vartheta>p$ and $t_{0}>0$ such that
    $0<\vartheta F(t)\leq tf(t),\quad\forall\,t>t_{0},\text{ where }F(t)=\int_{0}^{t}f(\tau)\,\mathrm{d}\tau.$
We consider the solution space for (SP) as
${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}:=\overline{{\mathcal{C}}_{c}^{\infty}({\mathbb{R}}^{N})}^{\left\lVert\cdot\right\rVert_{s,p}},\text{ where }\|u\|_{s,p}:=\bigg{(}\iint_{{\mathbb{R}}^{2N}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{N+sp}}\ {\mathrm{d}x\mathrm{d}y}\bigg{)}^{\frac{1}{p}}.$
The homogeneous fractional Sobolev space ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ has the following characterization (see [8, Theorem 3.1]):
${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}=\left\{u\in L^{\frac{Np}{N-sp}}({\mathbb{R}}^{N}):\left\lVert u\right\rVert_{s,p}<\infty\right\}.$ (1.2)
The fractional $p$-Laplace operator $(-\Delta)_{p}^{s}$ is defined as
$(-\Delta)_{p}^{s}u(x)=2\lim_{\epsilon\rightarrow 0^{+}}\int_{\mathbb{R}^{N}\backslash B_{\epsilon}(x)}\frac{|u(x)-u(y)|^{p-2}(u(x)-u(y))}{|x-y|^{N+sp}}\,{\mathrm{d}y},\quad\text{for}~{}x\in\mathbb{R}^{N},$
where $B_{\epsilon}(x)$ is the ball of radius $\epsilon$ and centred at $x$. A function $u\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ is called a weak solution to (SP) if it satisfies the following identity:
$\begin{split}\iint_{{\mathbb{R}}^{2N}}\frac{\lvert u(x)-u(y)\rvert^{p-2}(u(x)-u(y))(\phi(x)-\phi(y))}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}\\ =\int_{{\mathbb{R}}^{N}}g(x)f_{a}(u)\phi(x)\,{\rm d}x,\quad\forall\,\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}.\end{split}$
We call (SP) a semipositone problem since the term $g(x)f_{a}(u)$ is strictly negative on some part of the regions $\{u\leq 0\}$ and $\{u>0\}$ near $u=0$. Semipositone problems arise in mathematical biology, population dynamics, control theory, etc. (see [38, 16] and the references therein). Mathematicians have used several techniques to prove the existence of positive solution for semipositone problems, which include fixed point theory, sub and super-solution methods, and variational methods. These methods help to establish conditions under which solutions exist and (possibly under some additional hypotheses) provide insights into the qualitative behaviour of these solutions. Unlike positone problems, where the strong maximum principle guarantees the positivity of a non-negative solution, semipositone problems arise when the solution lives in regions where the source term is negative.
The study of semipositone problems began from the work [13] of Brown and Shivaji while studying the bifurcation theory for the perturbed problem $-\Delta u=\lambda(u-u^{3})-\epsilon$ in $\Omega$ and $u>0$ in $\Omega$, where $\lambda,\epsilon>0$ and $\Omega$ is a bounded domain. Subsequently, numerous authors have investigated the existence and the qualitative aspects of positive solutions to local semipositone problems across various domains. For bounded domain $\Omega$, relevant studies can be found in [21, 18, 15, 17, 19, 23, 3, 14, 36, 28, 22, 39, 34]. In the context where $\Omega$ is the exterior of a bounded domain, we refer [20, 37, 33, 32, 24] and the references therein. In [2], Alves et al. first studied the semipositone problem within the entire domain ${\mathbb{R}}^{N}$, as described by the equation $-\Delta u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N};N\geq 3$ with $f(0)=0$ and $a>0$. The function $f\in\mathcal{C}({\mathbb{R}}^{+})$ is locally Lipschitz, has subcritical and the Ambrosetti-Rabinowitz (A-R) type growth. The weight $g$ is positive and bounded by a radial function $P\in{\mathcal{C}}({\mathbb{R}}^{+})$, where $P$ satisfies $(a)\int_{{\mathbb{R}}^{N}}|x|^{2-N}P(|x|)\,{\rm d}x<\infty$, $(b)$ $P(\lvert\cdot\rvert)\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, and $(c)$ $|x|^{N-2}\int_{{\mathbb{R}}^{N}}P(|y|)|x-y|^{-N+2}\,{\mathrm{d}y}\leq C$, for all $x\in{\mathbb{R}}^{N}\setminus\{0\}$ and for some constant $C>0$. For $p\in(1,N)$, in [1], Santos et al. investigated the nonlinear variant $-\Delta_{p}u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N}$, where $f(0)=0,f\in{\mathcal{C}}({\mathbb{R}}^{+})$ exhibits subcritical with A-R type growth and the weight $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ satisfies $g(x)<B|x|^{-\vartheta}$ for $x\neq 0$, with $\vartheta>N$ and $B>0$. Meanwhile, the study in [6] focused on $\Delta^{2}u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N};N\geq 5$ where $f(0)=0$, $f\in{\mathcal{C}}({\mathbb{R}}^{+})$ satisfies weaker A-R type growth.
Generally, the techniques used to prove the existence of positive solution for semipositone problems defined on ${\mathbb{R}}^{N}$ differ from those defined on smaller domains. The authors in [2] studied an auxiliary problem $-\Delta u=g(x)f_{a}(u)$ in $\mathbb{R}^{N}$, where $f_{a}\in{\mathcal{C}}({\mathbb{R}})$ is defined as in (1.1). Subsequently, their main ideas to obtain positive solution are as follows: establish uniform boundedness of the weak solutions $\{u_{a}\}$ in $L^{\infty}({\mathbb{R}}^{N})$ (using the regularity estimate by Brezis and Kato in [12]), for $a$ near zero prove uniform convergence of $\{u_{a}\}$ (using the Riesz potential for the Laplace operator) to a non-negative function $\tilde{u}$ which is a weak solution of some positone problem, find the positivity of $\tilde{u}$ (applying the strong maximum principle), and finally (again using the Riesz potential along with the assumption $(c)$) obtain the positivity of $u_{a}$ for $a$ near zero. For $p\neq 2$, the Riesz representation for the $p$-Laplace operator is unavailable. The authors in [1] considered $-\Delta_{p}u=g(x)f_{a}(u)$, with a discontinuous function $f_{a}$ defined as $f_{a}(t)=f(t)-a$ for $t\geq 0$, $f_{a}(t)=0$ for $t<0$, and used a non-smooth variational approach. A key benefit of non-smooth analysis is that the critical point of the energy functional remains non-negative despite not being a weak solution to the problem. They established the existence of a positive critical point by constructing an explicit positive radial solution to a certain non-local equation and applying the comparison principle with that solution. This positive critical point ultimately serves as weak solution. The authors in [6] applied a similar technique as in [2] to obtain a positive solution.
Few articles are available in the literature dealing with non-local semipositone problems. In this direction, for $N>2s$ and $\Omega$ bounded, the authors in [25] studied $(-\Delta)^{s}u=\lambda(u^{q}-1)+\mu u^{r}$ in $\Omega$; $u=0$ in ${\mathbb{R}}^{N}\setminus\Omega$, where $\lambda,\mu>0$, $q\in(0,1)$ and $r\in(1,\frac{N+2s}{N-2s})$. Under certain lower bound of $\lambda$, they constructed a positive subsolution when $\mu=0$ and showed that there exists at least one positive solution when $0<\mu<\mu_{\lambda}$. For $N>p\geq 2$ and $\Omega$ bounded, the authors in [35] proved that $(-\Delta)_{p}^{s}u=\lambda f(u)$ in $\Omega$; $u=0$ in ${\mathbb{R}}^{N}\setminus\Omega$ admits positive solution, provided $\lambda>0$ is small. They used regularity of weak solution up to the boundary of $\Omega$ and Hopf’s Lemma for $(-\Delta)_{p}^{s}$. Recently, in [5], the author studies $(-\Delta)^{s}u=g(x)\left(f(u)-a\right)$ in $\mathbb{R}^{N}$ with $f(0)=0,f\in{\mathcal{C}}({\mathbb{R}}^{+})$ is locally Lipschitz, satisfies subcritical and weaker A-R type growth. Whereas, $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ satisfies $|x|^{N-2s}\int_{{\mathbb{R}}^{N}}g(y)|x-y|^{-N+2s}\,{\mathrm{d}y}\leq C(g),x\in{\mathbb{R}}^{N}\setminus\{0\},C(g)>0$. The existence of a positive solution is obtained employing similar techniques as in [2].
In this paper, we aim to study (SP), a non-local analogue of [1]. To our knowledge, non-linear non-local semipositone problems on the whole of $\mathbb{R}^{N}$ have not been addressed in the literature. The weight function $g$ falls within both $L^{1}(\mathbb{R}^{N})$ and $L^{\infty}(\mathbb{R}^{N})$ spaces, adheres to the bounds specified in (1.4). Meanwhile, the function $f$ meets subcritical and A-R type growth, as outlined in ($\bf f{1}$) and ($\bf f{2}$). Depending on the parameter $a$, our principal goal is to establish the existence of a positive solution to (SP). The nonsmooth variational technique as in [1] is not readily adapted for $s\in(0,1)$ due to the regularity constraints associated with the critical points of the non-smooth energy functional of (SP). This leads us to follow a different approach from [1]. We consider an energy functional associated with (SP), which has a ${\mathcal{C}}^{1}$ variational structure (see (2.1)). Applying the mountain pass theorem, we establish the existence of a mountain pass critical point for the energy functional, which corresponds to a mountain pass solution of (SP). The following theorem combines our main results.
Theorem 1.1.
Let $s\in(0,1),N\geq 1$, and $p\in(1,\frac{N}{s})$. Assume that $f$ satisfies ($\bf f{1}$) and ($\bf f{2}$). Let $g$ be a positive function with $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$. Then the following holds:
  1. (a)
    There exists $a_{1}>0$ such that for each $a\in(0,a_{1})$, (SP) admits a mountain pass solution $u_{a}$. Moreover, there exists a constant $C>0$ such that $\|u_{a}\|_{s,p}\leq C$ for all $a\in(0,a_{1})$.
  2. (b)
    In addition, we assume that $f$ satisfies the following condition at infinity:
    1. ($\bf\tilde{f{1}}$)
      $\lim\limits_{t\rightarrow\infty}\frac{f(t)}{t^{\frac{Np}{N-sp}-1}}=0$.
    Then for every $r\in[\frac{Np}{N-sp},\infty]$, there exists $C=C(r,N,s,p,f,g,a_{1})$ such that
    $\left\lVert u_{a}\right\rVert_{L^{r}({\mathbb{R}}^{N})}\leq C,\quad\forall\,a\in(0,a_{1}).$ (1.3)
  3. (c)
    Further, let $p>\frac{2N}{N+2s}$. Then there exists $\tilde{a}\in(0,a_{1})$ such that $u_{a}\geq 0$ a.e. in ${\mathbb{R}}^{N}$ for every $a\in(0,\tilde{a})$.
  4. (d)
    Furthermore, suppose $g$ satisfies the following bound:
    $g(x)\leq\frac{B}{|x|^{\beta(p-1)+sp}},\text{ for some constant }B>0\text{ and }x\neq 0,$ (1.4)
    where $\frac{N-sp}{p-1}<\beta<\frac{N}{p-1}$. Then, there exists $\hat{a}\in(0,\tilde{a})$ such that $u_{a}>0$ a.e. in ${\mathbb{R}}^{N}$ for every $a\in(0,\hat{a})$.
To show the positivity of the solution $u_{a}$, we first show that the sequence of solutions $\{u_{a}\}$ uniformly converges to a positive function in ${\mathcal{C}}({\mathbb{R}}^{N})$ as $a\rightarrow 0$ (see Proposition 3.5, Proposition 3.6, and Proposition 3.8). For $a$ near zero, we obtain an explicit positive subsolution of (SP) on the exterior of a ball in ${\mathbb{R}}^{N}$, following the approach in [11, Lemma 3.4]. Subsequently, we obtain the positivity of $u_{a}$ using the comparison principle [11, Theorem 2.7]. The strategy to prove Theorem 1.1-((c) and (d)) is different from [5, 2], where the Riesz potential for a linear operator plays a major role.
The rest of the paper is organized as follows. In Section 2, we set up a functional framework for (SP). Section 3 covers the existence and various qualitative properties of solutions to (SP). This section contains the proof of Theorem 1.1. In the Appendix, we provide some technical lemmas.
2. Functional frameworks for the problem
To obtain the existence of non-trivial solutions to (SP), this section studies variational settings. We fix some notations that will be used throughout this paper for brevity.
Notation: (i) We denote $X$ as a real Banach space endowed with the norm $\|\cdot\|_{X}$.
(ii) $X^{*}$ denotes the dual of $X$.
(iii) We denote $\|\cdot\|_{*}$ as the norm on $({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})})^{*}$.
(iv) For $p\in[1,\infty]$, the $L^{p}({\mathbb{R}}^{N})$ norm of a function $u$ is denoted as $\left\lVert u\right\rVert_{p}$.
(v) For $s\in(0,1)$ and $p\in(1,\frac{N}{s})$, $p_{s}^{*}=\frac{Np}{N-sp}$ is the non-local critical exponent.
(vi) We denote $\Phi:{\mathbb{R}}\rightarrow{\mathbb{R}}$ which is defined as $\Phi(t)=|t|^{p-2}t$.
(vii) We denote $C,\widetilde{C},C_{1},C_{2},C_{3}$ as positive constants.
(viii) $B_{r}$ denotes an open ball of radius $r$ with centre at origin.
(ix) For $A\subset{\mathbb{R}}^{N}$, $A^{c}$ denotes the complement of $A$, i.e., $A^{c}={\mathbb{R}}^{N}\setminus A$.
(x) For $s\in(0,1)$ and $p\in(1,\infty)$,
$L_{sp}^{p-1}({\mathbb{R}}^{N}):=\left\{u\in L_{\text{loc}}^{p-1}({\mathbb{R}}^{N}):\int_{{\mathbb{R}}^{N}}\frac{\lvert u(x)\rvert^{p-1}}{(1+\lvert x\rvert)^{N+sp}}\,{\rm d}x<\infty\right\}.$
(xi) For $s\in(0,1)$ and $p\in(1,\infty)$,
$W^{s,p}(\Omega):=\left\{u\in L^{p}(\Omega):\lvert u\rvert^{p}_{W^{s,p}(\Omega)}:=\iint_{\Omega\times\Omega}\frac{\lvert u(x)-u(y)\rvert^{p}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}<\infty\right\},$
and
$W^{s,p}_{\text{loc}}(\Omega):=\left\{u\in L^{p}_{\text{loc}}(\Omega):u\in W^{s,p}(\Omega_{1}),\text{ for any relatively compact open set }\Omega_{1}\subset\Omega\right\}.$
For $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ and $a\geq 0$, we define the maps $K_{a}$ and $I_{a}$ on ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ as follows
$K_{a}(u):=\int_{{\mathbb{R}}^{N}}g(x)F_{a}(u(x))\,{\rm d}x,\text{ and }I_{a}(u):=\frac{1}{p}\|u\|_{s,p}^{p}-K_{a}(u).$
Notice that $K_{a},I_{a}\in{\mathcal{C}}^{1}({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})},{\mathbb{R}})$ and the corresponding Fréchet derivatives are given by
$\begin{split}&K_{a}^{\prime}(u)(v)=\int_{{\mathbb{R}}^{N}}g(x)f_{a}(u)v\,{\rm d}x,\text{ and }\\ &I_{a}^{\prime}(u)(v)=\iint_{{\mathbb{R}}^{2N}}\frac{\Phi(u(x)-u(y))(v(x)-v(y))}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}-K_{a}^{\prime}(u)(v),\;\forall\,v\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}.\end{split}$ (2.1)
Moreover, every critical point of $I_{a}$ corresponds to a solution of (SP). Before discussing further properties of $K_{a},I_{a}$, we identify the upper and lower bounds of $f,f_{a},$ and their primitives. Clearly, the function $f_{a}\in{\mathcal{C}}({\mathbb{R}})$ and its primitive $F_{a}$ is defined as $F_{a}(t)=F(t)-at$ for $t\geq 0$, $F_{a}(t)=-\frac{at^{2}}{2}-at$ for $t\in[-1,0]$ and $F_{a}(t)=\frac{a}{2}$ for $t\leq-1$. From ($\bf f{1}$),
$\lim\limits_{t\rightarrow 0^{+}}\frac{f(t)}{t^{p-1}}=0\Rightarrow\text{ for every }\epsilon>0,\text{ there exists }t_{1}(\epsilon)>0\text{ suchthat }f(t)<\epsilon t^{p-1}\text{ for }t\in(0,t_{1}(\epsilon)).$
$\lim\limits_{t\rightarrow\infty}\frac{f(t)}{t^{\gamma-1}}\leq C(f)\Rightarrow f(t)\leq C(f,t_{1}(\epsilon))t^{\gamma-1}\text{ for }t\geq t_{1}(\epsilon).$
Hence, we have the following bounds for $\gamma\in(p,p^{*}_{s}]$:
$|f_{a}(t)|\leq\epsilon|t|^{p-1}+C(f,t_{1}(\epsilon))|t|^{\gamma-1}+a\text{ and }|F_{a}(t)|\leq\epsilon|t|^{p}+C(f,t_{1}(\epsilon))|t|^{\gamma}+a|t|\text{ for }t\in{\mathbb{R}}.$ (2.2)
Again using the subcritical growth on $f$, $f(t)\leq C(f)t^{\gamma-1}$, for $t>t_{2}$, for some $t_{2}>0$. The continuity of $f$ infers that $f(t)\leq C$ on $[0,t_{2}]$. Hence for $a\in(0,\tilde{a})$, we obtain
$|f_{a}(t)|\leq C(1+|t|^{\gamma-1})\text{ and }|F_{a}(t)|\leq C(|t|+|t|^{\gamma})\text{ for }t\in{\mathbb{R}},\text{ where }C=C(f,t_{2},\tilde{a}).$ (2.3)
By ($\bf f{2}$) and the continuity of $F(t)$, there exist $M_{1},M_{2}>0$ such that
$F(t)\geq M_{1}t^{\vartheta}-M_{2},\;\forall\,t\geq 0.$ (2.4)
Remark 2.1 (A-R condition of $f_{a}$).
For $t>t_{0}$, it follows from ($\bf f{2}$) that
$\vartheta F_{a}(t)=\vartheta F(t)-\vartheta at\leq tf(t)-at=tf_{a}(t).$
For $t\in[0,t_{0}]$, by continuity of $F$, there exists $M>0$ independent of $a$, such that
$\vartheta F_{a}(t)=\vartheta F(t)-\vartheta at\leq M-at\leq M-at+tf(t)=tf_{a}(t)+M.$
For $t\in[-1,0]$ and $a\in(0,\tilde{a})$, observe that $\vartheta F_{a}(t)\leq-\vartheta at-\frac{at^{2}}{2}$ and $tf_{a}(t)=-at^{2}-at$. Then we have the following estimate:
$\vartheta F_{a}(t)-tf_{a}(t)\leq-(\vartheta-1)at+\frac{at^{2}}{2}\leq-(\vartheta-1)\tilde{a}t+\frac{\tilde{a}}{2}\leq\vartheta\tilde{a}.$
For $t\leq-1$, we have
$\vartheta F_{a}(t)=\frac{\vartheta a}{2}\leq\frac{\vartheta\tilde{a}}{2}\leq tf_{a}(t)+\vartheta\tilde{a}.$
By choosing $M_{3}=\max\{M,\vartheta\tilde{a}\}$, we obtain the following Ambrosetti-Rabinowitz (A-R) condition for $f_{a}$:
$\vartheta F_{a}(t)\leq tf_{a}(t)+M_{3},\;\forall\,t\in{\mathbb{R}}\text{ and }a\in(0,\tilde{a}),$ (2.5)
where $M_{3}$ is independent of $a$ and $t$.
The following proposition states some compact embeddings of the solution space for (SP) into the spaces of locally integrable functions and weighted Lebesgue spaces.
Proposition 2.2.
Let $p\in(1,\frac{N}{s})$ and $q\in[1,p_{s}^{*})$. Then the following hold:
  1. (i)
    The embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L_{\text{loc}}^{q}({\mathbb{R}}^{N})$ is compact.
  2. (ii)
    Suppose $g\in L^{\alpha}({\mathbb{R}}^{N})$ for $\alpha=\frac{p_{s}^{*}}{p_{s}^{*}-q}$. Then the embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{q}({\mathbb{R}}^{N},|g|)$ is compact.
Proof.
(i) Proof follows using [9, Lemma A.1] and the same arguments as given in [5, Proposition 2.1].
(ii) Assume that $u_{n}\rightharpoonup u$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. We need to prove $u_{n}\rightarrow u$ in $L^{q}({\mathbb{R}}^{N},|g|)$. The space ${\mathcal{C}}_{c}({\mathbb{R}}^{N})$ is dense in $L^{\alpha}({\mathbb{R}}^{N})$ and hence for every $\epsilon>0$ we take $g_{\epsilon}\in{\mathcal{C}}_{c}({\mathbb{R}}^{N})$ with $K:=\mathrm{supp}(g_{\epsilon})$ such that
$\|g-g_{\epsilon}\|_{\alpha}<\frac{\epsilon}{2L},$
where $L:=\sup\{\|u_{n}-u\|_{p_{s}^{*}}^{q}:n\in{\mathbb{N}}\}$. Now, using the triangle and Hölder inequalities, we obtain the following estimates:
$\int_{{\mathbb{R}}^{N}}|g||u_{n}-u|^{q}\,{\rm d}x$ $\leq\int_{{\mathbb{R}}^{N}}|g-g_{\epsilon}||u_{n}-u|^{q}\,{\rm d}x+\int_{{\mathbb{R}}^{N}}|g_{\epsilon}||u_{n}-u|^{q}\,{\rm d}x$
$\leq\|g-g_{\epsilon}\|_{\alpha}\|u_{n}-u\|_{p_{s}^{*}}^{q}+\int_{K}|g_{\epsilon}||u_{n}-u|^{q}\,{\rm d}x\leq\frac{\epsilon}{2}+M\int_{K}|u_{n}-u|^{q}\,{\rm d}x.$ (2.6)
The above constant $M$ is the upper bound of $g_{\epsilon}$ on the compact set $K$. Moreover, the embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{q}_{\text{loc}}({\mathbb{R}}^{N})$ is compact. Therefore, there exists $n_{1}\in{\mathbb{N}}$ such that up to a subsequence
$\int_{K}|u_{n}-u|^{q}\,{\rm d}x<\frac{\epsilon}{2M},\;\forall\,n\geq n_{1}.$ (2.7)
From (2) and (2.7), we deduce that
$\int_{{\mathbb{R}}^{N}}|g||u_{n}-u|^{q}\,{\rm d}x<\epsilon,\;\forall\,n\geq n_{1}.$
Since $\epsilon>0$ is arbitrary, we get the desired result. ∎
Now, we prove the compactness of $K_{a}$ using a similar splitting argument as above.
Proposition 2.3.
Let $p\in(1,\frac{N}{s})$. Assume that $f$ satisfies ($\bf f{1}$) and $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$. Then $K_{a}$ is compact on ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ for every $a\geq 0$.
Proof.
Let $u_{n}\rightharpoonup u$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Since ${\mathcal{C}}_{c}({\mathbb{R}}^{N})$ is dense in both $L^{\frac{p_{s}^{*}}{p_{s}^{*}-1}}({\mathbb{R}}^{N})$ and $L^{\frac{p_{s}^{*}}{p_{s}^{*}-\gamma}}({\mathbb{R}}^{N})$, for every given $\epsilon>0$ we take $g_{\epsilon}\in{\mathcal{C}}_{c}({\mathbb{R}}^{N})$ such that
$\|g-g_{\epsilon}\|_{\frac{p_{s}^{*}}{p_{s}^{*}-1}}+\|g-g_{\epsilon}\|_{\frac{p_{s}^{*}}{p_{s}^{*}-\gamma}}<\frac{\epsilon}{L},$ (2.8)
where $L:=\sup\{\|u_{n}\|_{p_{s}^{*}}+\|u\|_{p_{s}^{*}}+\|u_{n}\|_{p_{s}^{*}}^{\gamma}+\|u\|_{p_{s}^{*}}^{\gamma}:n\in{\mathbb{N}}\}$. Since the sequence $\{\|u_{n}\|_{s,p}\}$ is bounded in $\mathbb{R}^{+}$, using the continuous embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p_{s}^{*}}({\mathbb{R}}^{N})$ we see that $\{\|u_{n}\|_{p_{s}^{*}}\}$ is also bounded in $\mathbb{R}^{+}$. Therefore, $L<\infty$. For $a\geq 0$, we write
$\big{|}K_{a}(u_{n})-K_{a}(u)\big{|}\leq\int_{{\mathbb{R}}^{N}}g(x)\left|\big{(}F_{a}(u_{n}(x))-F_{a}(u(x))\big{)}\right|\,{\rm d}x$
$\leq\int_{{\mathbb{R}}^{N}}|g-g_{\epsilon}|\big{|}F_{a}(u_{n}(x))-F_{a}(u(x))\big{|}\,{\rm d}x+\int_{{\mathbb{R}}^{N}}|g_{\epsilon}|\big{|}F_{a}(u_{n}(x))-F_{a}(u(x))\big{|}\,{\rm d}x:=\text{I}+\text{II}.$ (2.9)
Now using (2.3) and Hölder’s inequality with (2.8), we estimate the first integral as
I $\leq\int_{{\mathbb{R}}^{N}}|g-g_{\epsilon}|\big{(}|F_{a}(u_{n})|+|F_{a}(u)|\big{)}\,{\rm d}x\leq C\int_{{\mathbb{R}}^{N}}|g-g_{\epsilon}|\big{(}|u_{n}|+|u_{n}|^{\gamma}+|u|+|u|^{\gamma}\big{)}\,{\rm d}x$
$\leq C\big{(}\|g-g_{\epsilon}\|_{\frac{p_{s}^{*}}{p_{s}^{*}-1}}(\|u_{n}\|_{p_{s}^{*}}+\|u\|_{p_{s}^{*}})+\|g-g_{\epsilon}\|_{\frac{p_{s}^{*}}{p_{s}^{*}-\gamma}}(\|u_{n}\|_{p_{s}^{*}}^{\gamma}+\|u\|_{p_{s}^{*}}^{\gamma})\big{)}<C\epsilon,$ (2.10)
where $C=C(f,a)$. Further, we estimate the following integral
$\text{II}=\int_{{\mathbb{R}}^{N}}|g_{\epsilon}|\big{|}F_{a}(u_{n})-F_{a}(u)\big{|}\,{\rm d}x\leq M\int_{K}\big{|}F_{a}(u_{n})-F_{a}(u)\big{|}\,{\rm d}x,$ (2.11)
where $K$ is the support of $g_{\epsilon}$ and $M=\|g_{\epsilon}\|_{\infty}$. Since ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ is compactly embedded into $L^{\gamma}_{\text{loc}}({\mathbb{R}}^{N})$ (Proposition 2.2), up to a subsequence we get $u_{n}\rightarrow u$ in $L^{\gamma}(K)$, and subsequently $u_{n}(x)\rightarrow u(x)$ a.e. in $K$. Moreover, $|F_{a}(u_{n})|\leq C(|u_{n}|+|u_{n}|^{\gamma})$ and $\int_{K}|u_{n}|\,{\rm d}x\rightarrow\int_{K}|u|\,{\rm d}x,~{}~{}\int_{K}|u_{n}|^{\gamma}\,{\rm d}x\rightarrow\int_{K}|u|^{\gamma}\,{\rm d}x$. Therefore, the generalized dominated convergence theorem yields $F_{a}(u_{n})\rightarrow F_{a}(u)$ in $L^{1}(K)$. Thus from (2.11),
$\int_{{\mathbb{R}}^{N}}|g_{\epsilon}|\big{|}F_{a}(u_{n})-F_{a}(u)\big{|}\,{\rm d}x\rightarrow 0~{}~{}\text{ as }n\rightarrow\infty.$
Now we conclude from (2) and (2) that $K_{a}(u_{n})\rightarrow K_{a}(u)$ as $n\rightarrow\infty$. ∎
Now, depending on the values of $a$, we verify the mountain pass geometry for the energy functional $I_{a}$.
Lemma 2.4.
Let $p\in(1,\frac{N}{s})$. Let $f$ satisfies ($\bf f{1}$), ($\bf f{2}$) and $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ be positive. Then the following hold:
  • (i)
    There exists $\beta,\delta>0$, and $a_{1}>0$ such that $I_{a}(u)\geq\delta$ for $\left\lVert u\right\rVert_{s,p}=\beta$ whenever $a\in(0,a_{1})$.
  • (ii)
    There exists $v\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ with $\|v\|_{s,p}>\beta$ such that $I_{a}(v)<0$, for all $a>0$.
Proof.
(i) The functional $I_{a}:{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\rightarrow{\mathbb{R}}$ is given by
$I_{a}(u)=\frac{1}{p}\iint_{{\mathbb{R}}^{2N}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{N+sp}}\ {\mathrm{d}x\mathrm{d}y}-\int_{{\mathbb{R}}^{N}}g(x)F_{a}(u)\,{\rm d}x.$
Using (2.2), we estimate
$\int_{{\mathbb{R}}^{N}}g(x)F_{a}(u)\,{\rm d}x$ $\leq\int_{{\mathbb{R}}^{N}}g(x)\big{(}\epsilon|u|^{p}+C(f,\epsilon)|u|^{\gamma}+a|u|\big{)}\,{\rm d}x$
$=\epsilon\int_{{\mathbb{R}}^{N}}g(x)|u|^{p}\,{\rm d}x+C(f,\epsilon)\int_{{\mathbb{R}}^{N}}g(x)|u|^{\gamma}\,{\rm d}x+a\int_{{\mathbb{R}}^{N}}g(x)|u|\,{\rm d}x$
$\leq\epsilon C_{1}\left\lVert u\right\rVert_{s,p}^{p}+C(f,\epsilon)C_{2}\left\lVert u\right\rVert_{s,p}^{\gamma}+aC_{3}\|u\|_{s,p},$
where $C_{1},C_{2}$ and $C_{3}$ are the embedding constants of ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{q}({\mathbb{R}}^{N},g);q\in[1,p^{*}_{s})$ (Proposition 2.2). In particular, for $\left\lVert u\right\rVert_{s,p}=\beta$,
$I_{a}(u)$ $\geq\frac{1}{p}\beta^{p}-\epsilon C_{1}\beta^{p}-C(f,\epsilon)C_{2}\beta^{\gamma}-aC_{3}\beta$
$=\beta^{p}\bigg{(}\frac{1}{p}-\epsilon C_{1}-C(f,\epsilon)C_{2}\beta^{\gamma-p}\bigg{)}-aC_{3}\beta.$ (2.12)
We choose $\epsilon<(pC_{1})^{-1}$. Then we write $I_{a}(u)\geq A(\beta)-aC_{3}\beta$, where $A(\beta)=C\beta^{p}(1-\widetilde{C}\beta^{\gamma-p})$ with $C,\widetilde{C}$ independent of $a$. Let $\beta_{1}$ be the first non-trivial zero of $A$. For $\beta<\beta_{1}$, we fix $a_{1}\in(0,\frac{A(\beta)}{C_{3}\beta})$ and $\delta=A(\beta)-a_{1}C_{3}\beta$. Thus, by (2) we get $I_{a}(u)\geq\delta$ for every $a\in(0,a_{1})$.
(ii) Let $\varphi\in C_{c}^{\infty}({\mathbb{R}}^{N})\setminus\{0\},\varphi\geq 0$ and $\|\varphi\|_{s,p}=1$. For $t\geq 0$, we have
$I_{a}(t\varphi)=\frac{t^{p}}{p}\iint_{{\mathbb{R}}^{2N}}\frac{|\varphi(x)-\varphi(y)|^{p}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}-\int_{{\mathbb{R}}^{N}}g(x)\left(F(t\varphi)-at\phi\right)\,{\rm d}x.$
Now using the A-R condition (2.4) of $F$, we obtain the following
$I_{a}(t\varphi)$ $\leq\frac{t^{p}}{p}\|\varphi\|_{s,p}^{p}-M_{1}t^{\vartheta}\int_{{\mathbb{R}}^{N}}g(x)(\varphi(x))^{\vartheta}\,{\rm d}x+M_{2}\int_{{\mathbb{R}}^{N}}g(x)\,{\rm d}x+at\int_{{\mathbb{R}}^{N}}g(x)\varphi(x)\,{\rm d}x$
$\leq\frac{t^{p}}{p}-M_{1}t^{\vartheta}\int_{{\mathbb{R}}^{N}}g(x)(\varphi(x))^{\vartheta}\,{\rm d}x+M_{2}\|g\|_{1}+at\left\lVert\varphi\right\rVert_{\infty}\left\lVert g\right\rVert_{1}.$ (2.13)
Using the fact $\vartheta>p>1$, it is easy to see that $I_{a}(t\varphi)\rightarrow-\infty$ as $t\rightarrow+\infty$. Thus, there exists a $t_{1}>\beta$ such that $I_{a}(t\varphi)<0$ for $t>t_{1}$. Therefore, the required function is $v=t\varphi$ with $t>t_{1}$. ∎
Definition 2.5 (Palais Smale condition).
Let $J:{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\rightarrow{\mathbb{R}}$ be a continuously differentiable functional. Then $J$ satisfies the Palais Smale (PS) condition if every sequence $\{u_{n}\}\subset{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ with $\{J(u_{n})\}$ is bounded in ${\mathbb{R}}$ and $J^{\prime}(u_{n})\rightarrow 0$ in $({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})})^{*}$ possesses a convergent subsequence.
Proposition 2.6.
Let $p\in(1,\frac{N}{s})$. Let $f$ satisfies ($\bf f{1}$), ($\bf f{2}$), and $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ be positive. Then $I_{a}$ satisfies the (PS) condition for every $a\geq 0$.
Proof.
Let $\{u_{n}\}$ be a sequence in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ such that $\{I_{a}(u_{n})\}$ is bounded in ${\mathbb{R}}$ and $I_{a}^{\prime}(u_{n})\rightarrow 0$ in $({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})})^{*}$. We need to show that the sequence $\{u_{n}\}$ has a strongly convergent subsequence in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. First, we show that $\{u_{n}\}$ is a bounded sequence in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Since $\{I_{a}(u_{n})\}$ is bounded, we have $|I_{a}(u_{n})|\leq M$ for some $M>0$, which further implies
$\frac{1}{p}\left\lVert u_{n}\right\rVert_{s,p}^{p}-\int_{{\mathbb{R}}^{N}}g(x)F_{a}(u_{n})\,{\rm d}x\leq M,\;\forall\,n\in{\mathbb{N}}.$ (2.14)
Next, using $I_{a}^{\prime}(u_{n})\rightarrow 0$ in $({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})})^{*}$, there exists $n_{1}\in{\mathbb{N}}$ such that for every $n\geq n_{1}$, we have $\lvert I_{a}^{\prime}(u_{n})(u_{n})\rvert\leq\left\lVert u_{n}\right\rVert_{s,p}$. Thus we obtain
$-\left\lVert u_{n}\right\rVert_{s,p}-\left\lVert u_{n}\right\rVert_{s,p}^{p}\leq-\int_{{\mathbb{R}}^{N}}g(x)f_{a}(u_{n})u_{n}\,{\rm d}x,\;\forall\,n\geq n_{1}.$ (2.15)
From (2.5) and (2.14), we see that
$\frac{1}{p}\left\lVert u_{n}\right\rVert_{s,p}^{p}-\frac{1}{\vartheta}\int_{{\mathbb{R}}^{N}}g(x)f_{a}(u_{n})u_{n}\,{\rm d}x-\frac{1}{\vartheta}M_{3}\left\lVert g\right\rVert_{1}\leq M,$ (2.16)
and further by (2.15) and (2.16),
$\bigg{(}\frac{1}{p}-\frac{1}{\vartheta}\bigg{)}\left\lVert u_{n}\right\rVert_{s,p}^{p}-\frac{1}{\vartheta}\left\lVert u_{n}\right\rVert_{s,p}\leq M+\frac{1}{\vartheta}M_{3}\left\lVert g\right\rVert_{1},\;\forall\,n\geq n_{1}.$
The above inequality infers that $\{u_{n}\}$ is bounded in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. By reflexivity of ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, there exists $u\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ such that up to a subsequence $u_{n}\rightharpoonup u$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Now we consider the functional $J(v)=\frac{1}{p}\|v\|_{s,p}^{p}$ for $v\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Clearly, $J\in{\mathcal{C}}^{1}({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})},{\mathbb{R}})$. From (2.1), we write $J^{\prime}(u_{n})(u_{n}-u)=I_{a}^{\prime}(u_{n})(u_{n}-u)+K_{a}^{\prime}(u_{n})(u_{n}-u)$. Now we claim that $J^{\prime}(u_{n})(u_{n}-u)\rightarrow 0$ as $n\rightarrow\infty$. First, we prove $I_{a}^{\prime}(u_{n})(u_{n}-u)\rightarrow 0$ as $n\rightarrow\infty$. Recall that $I_{a}^{\prime}(u_{n})\rightarrow 0\text{ in }({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})})^{*}$ and $\{u_{n}\}$ is bounded in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Consequently, we deduce $|I_{a}^{\prime}(u_{n})(u_{n}-u)|\leq\|I_{a}^{\prime}(u_{n})\|_{*}\|u_{n}-u\|_{s,p}\rightarrow 0$ as $n\rightarrow\infty$. Next, we prove $K_{a}^{\prime}(u_{n})(u_{n}-u)\rightarrow 0$. Observe from (2.3) that
$|K_{a}^{\prime}(u_{n})(u_{n}-u)|$ $\leq\int_{{\mathbb{R}}^{N}}g(x)|f_{a}(u_{n})||u_{n}-u|\,{\rm d}x$
$\leq C(f,a)\int_{{\mathbb{R}}^{N}}g(x)\left(1+|u_{n}|^{\gamma-1}\right)|u_{n}-u|\,{\rm d}x.$ (2.17)
The sequence $\{|u_{n}|^{\gamma-1}\}$ is bounded in $L^{\frac{p_{s}^{*}}{\gamma-1}}({\mathbb{R}}^{N})$ (as $\{u_{n}\}$ is bounded in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$). Further,
$\gamma<p_{s}^{*}\Longleftrightarrow\frac{p_{s}^{*}}{p_{s}^{*}-(\gamma-1)}<p_{s}^{*}.$
Therefore, applying Proposition 2.2 and $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, we get $u_{n}\rightarrow u$ in $L^{\frac{p_{s}^{*}}{p_{s}^{*}-(\gamma-1)}}({\mathbb{R}}^{N},g)$. Thus, using Hölder’s inequality with conjugate pair $(\frac{p_{s}^{*}}{\gamma-1},\frac{p_{s}^{*}}{p_{s}^{*}-(\gamma-1)})$,
$\int_{{\mathbb{R}}^{N}}g(x)|u_{n}-u||u_{n}|^{\gamma-1}\,{\rm d}x\leq\left\lVert g\right\rVert_{\infty}^{\frac{\gamma-1}{p^{*}_{s}}}\left\lVert u_{n}-u\right\rVert_{L^{\frac{p_{s}^{*}}{p_{s}^{*}-(\gamma-1)}}({\mathbb{R}}^{N},g)}\||u_{n}|^{\gamma-1}\|_{\frac{p_{s}^{*}}{\gamma-1}}\rightarrow 0\text{ as}n\rightarrow\infty.$ (2.18)
Also, $\int_{{\mathbb{R}}^{N}}g(x)|u_{n}-u|\,{\rm d}x\rightarrow 0$ as $n\rightarrow\infty$ (by Proposition 2.2). Thus we have
$\int_{{\mathbb{R}}^{N}}g(x)\left(1+|u_{n}|^{\gamma-1}\right)|u_{n}-u|\,{\rm d}x\rightarrow 0\text{ as }n\rightarrow\infty.$ (2.19)
We infer from (2) and (2.19) that $K_{a}^{\prime}(u_{n})(u_{n}-u)\rightarrow 0$ as $n\rightarrow\infty$. Thus, our claim holds true. Since $J\in{\mathcal{C}}^{1}({{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})},{\mathbb{R}})$ and $u_{n}\rightharpoonup u$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we also get $J^{\prime}(u)(u_{n}-u)\rightarrow 0$ as $n\rightarrow\infty$. Further, we estimate the following using Hölder inequality:
$J^{\prime}(u_{n})(u_{n}-u)-J^{\prime}(u)(u_{n}-u)$
$~{}~{}=\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u_{n}(x)-u_{n}(y)\big{)}-\Phi\big{(}u(x)-u(y)\big{)}}{|x-y|^{N+sp}}\big{(}(u_{n}-u)(x)-(u_{n}-u)(y)\big{)}\,{\mathrm{d}x\mathrm{d}y}$
$~{}~{}=\iint_{{\mathbb{R}}^{2N}}\frac{|u_{n}(x)-u_{n}(y)|^{p}}{|x-y|^{N+sp}}\ {\mathrm{d}x\mathrm{d}y}+\iint_{{\mathbb{R}}^{2N}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{N+sp}}\ {\mathrm{d}x\mathrm{d}y}$
$~{}~{}\;-\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u_{n}(x)-u_{n}(y)\big{)}\big{(}u(x)-u(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}-\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u(x)-u(y)\big{)}\big{(}u_{n}(x)-u_{n}(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$~{}~{}\geq\left\lVert u_{n}\right\rVert_{s,p}^{p}+\left\lVert u\right\rVert_{s,p}^{p}-\left\lVert u_{n}\right\rVert_{s,p}^{p-1}\left\lVert u\right\rVert_{s,p}-\left\lVert u\right\rVert_{s,p}^{p-1}\left\lVert u_{n}\right\rVert_{s,p}$
$~{}~{}=\big{(}\left\lVert u_{n}\right\rVert_{s,p}^{p-1}-\left\lVert u\right\rVert_{s,p}^{p-1}\big{)}\big{(}\left\lVert u_{n}\right\rVert_{s,p}-\left\lVert u\right\rVert_{s,p}\big{)}\geq 0.$ (2.20)
Therefore, taking limit as $n\rightarrow\infty$ in (2), we get $\left\lVert u_{n}\right\rVert_{s,p}\rightarrow\left\lVert u\right\rVert_{s,p}$. Since ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ is uniformly convex Banach space, $u_{n}\rightarrow u$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. ∎
3. Existence and qualitative properties of the mountain pass solutions
The following theorem states the existence and uniform boundedness of the mountain pass solutions of (SP).
Theorem 3.1.
Let $p\in(1,\frac{N}{s})$. Assume that $f$ satisfies ($\bf f{1}$) and ($\bf f{2}$). Let $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ be positive. Let $a_{1}>0$ be given in Lemma 2.4. Then for each $a\in(0,a_{1})$, (SP) admits a mountain pass solution $u_{a}$. Moreover, there exists $C>0$ such that $\|u_{a}\|_{s,p}\leq C$ for all $a\in(0,a_{1})$.
Proof.
Consider $a_{1},\delta,v$ as given in Lemma 2.4. For $a\in(0,a_{1})$, using Lemma 2.4 and Proposition 2.6, we observe that all the hypotheses of the mountain pass theorem [4, Theorem 2.1] are verified. Therefore, by [4, Theorem 2.1], there exists a non-trivial critical point $u_{a}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ of $I_{a}$ satisfying
$I_{a}(u_{a})=c_{a}=\inf\limits_{\gamma\in\Gamma_{v}}\max\limits_{t\in[0,1]}I_{a}(\gamma(t))\geq\delta\text{ and }I_{a}^{\prime}(u_{a})=0,$ (3.1)
where $\Gamma_{v}:=\{\gamma\in C([0,1],{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}):\gamma(0)=0\text{ and }\gamma(1)=v\}$ and $c_{a}$ is the mountian pass level associated with $I_{a}$. Thus, $u_{a}$ is a non-trivial solution to (SP). To prove the uniform boundedness of $u_{a}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we first show that the set $\{I_{a}(u_{a}):a\in(0,a_{1})\}$ is uniformly bounded. We define a path $\tilde{\gamma}:[0,1]\rightarrow{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ by $\tilde{\gamma}(\sigma)=\sigma v$, where $v=t\varphi$ for some $t>t_{1}$ (for $t_{1}$ as in Lemma 2.4-(ii)), $\varphi\in{\mathcal{C}}_{c}^{\infty}({\mathbb{R}}^{N})\setminus\{0\},~{}\varphi\geq 0$ and $\left\lVert\varphi\right\rVert_{s,p}=1$. We see that $\tilde{\gamma}\in\Gamma_{v}$ because $\tilde{\gamma}(0)=0$ and $\tilde{\gamma}(1)=v$. Now from (3.1), we have
$I_{a}(u_{a})=c_{a}=\inf\limits_{\gamma\in\Gamma_{v}}\max\limits_{t\in[0,1]}I_{a}(\gamma(t))\leq\max\limits_{\sigma\in[0,1]}I_{a}(\tilde{\gamma}(\sigma))=\max\limits_{\sigma\in[0,1]}I_{a}(\sigma t\varphi).$ (3.2)
Now estimating as in (2), we obtain
$I_{a}(\sigma t\varphi)$ $\leq\frac{{\sigma}^{p}t^{p}}{p}\left\lVert\varphi\right\rVert_{s,p}^{p}-M_{1}{\sigma}^{\vartheta}t^{\vartheta}\int_{{\mathbb{R}}^{N}}g(x)(\varphi(x))^{\vartheta}\,{\rm d}x+M_{2}\int_{{\mathbb{R}}^{N}}g(x)\,{\rm d}x+a{\sigma}t\int_{{\mathbb{R}}^{N}}g(x)\varphi(x)\,{\rm d}x$
$\leq\frac{t^{p}}{p}+M_{2}\|g\|_{1}+a_{1}tC_{1}\left\lVert\varphi\right\rVert_{s,p}.$ (3.3)
From (3.2) and (3), there exists $C=C(N,s,p,M_{2},g,a_{1})$ such that
$I_{a}(u_{a})\leq C,\text{ for all }a\in(0,a_{1}).$ (3.4)
Using the uniform boundedness of $I_{a}(u_{a})$, we will show that the solutions $\{u_{a}\}$ are uniformly bounded in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. From (3.1), we write
$\left\lVert u_{a}\right\rVert_{s,p}^{p}-\int_{{\mathbb{R}}^{N}}g(x)f_{a}(u_{a})u_{a}\,{\rm d}x=0.$ (3.5)
Also, by (3.4) we have
$\frac{1}{p}\left\lVert u_{a}\right\rVert_{s,p}^{p}-\int_{{\mathbb{R}}^{N}}g(x)F_{a}(u_{a})\,{\rm d}x\leq C.$ (3.6)
Now first multiplying (3.5) by $\frac{1}{\vartheta}$ and then subtracting into (3.6) gives the following
$\bigg{(}\frac{1}{p}-\frac{1}{\vartheta}\bigg{)}\left\lVert u_{a}\right\rVert_{s,p}^{p}+\int_{{\mathbb{R}}^{N}}g(x)\bigg{(}\frac{1}{\vartheta}f_{a}(u_{a})u_{a}-F_{a}(u_{a})\bigg{)}\,{\rm d}x\leq C,$
and combining the above with (2.5) yields
$\bigg{(}\frac{1}{p}-\frac{1}{\vartheta}\bigg{)}\left\lVert u_{a}\right\rVert_{s,p}^{p}-\frac{1}{\vartheta}M_{3}\|g\|_{1}\leq C.$
Thus, there exists $C=C(N,s,p,f,g,a_{1})$ such that $\left\lVert u_{a}\right\rVert_{s,p}\leq C$ for every $a\in(0,a_{1})$. ∎
From Theorem 3.1 and the embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$ it is clear that $\{u_{a}:a\in(0,a_{1})\}$ is uniformly bounded on $L^{p^{*}_{s}}({\mathbb{R}}^{N})$. In the following proposition, using the Moser iteration technique, we discuss the uniform boundedness of $\{u_{a}:a\in(0,a_{1})\}$ over $L^{r}({\mathbb{R}}^{N})$ for every $r\in[p^{*}_{s},\infty]$.
Remark 3.2.
For $h\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we check that $h^{\pm}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ and $\left\lVert h^{\pm}\right\rVert_{s,p}\leq\left\lVert h\right\rVert_{s,p}$. We verify this for the positive part of $h$; a similar argument holds for the negative part. Set $A_{h}:=\{x\in{\mathbb{R}}^{N}:h(x)\geq 0\}$. Then we see that
$\left\lVert h^{+}\right\rVert_{s,p}^{p}=$ $\iint_{{\mathbb{R}}^{2N}}\frac{\lvert h^{+}(x)-h^{+}(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\left(\iint_{A_{h}\times A_{h}}+2\iint_{A_{h}\times(A_{h})^{c}}\right)\frac{\lvert h^{+}(x)-h^{+}(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$=\iint_{A_{h}\times A_{h}}\frac{\lvert h(x)-h(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}+2\iint_{A_{h}\times(A_{h})^{c}}\frac{\lvert h(x)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$\leq\left(\iint_{A_{h}\times A_{h}}+2\iint_{A_{h}\times(A_{h})^{c}}+\iint_{(A_{h})^{c}\times(A_{h})^{c}}\right)\frac{\lvert h(x)-h(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\left\lVert h\right\rVert_{s,p}^{p}.$
Proposition 3.3.
Let $p\in(1,\frac{N}{s})$. Let $f$ satisfies ($\bf f{1}$), ($\bf f{2}$), and $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ be positive. In addition, we assume that $f$ satisfies the following condition at infinity:
  1. ($\bf\tilde{f{1}}$)
    $\lim\limits_{t\rightarrow\infty}\frac{f(t)}{t^{p^{*}_{s}-1}}=0$.
Then for every $r\in[p^{*}_{s},\infty]$ there exists $C=C(r,N,s,p,f,g,a_{1})>0$ such that
$\left\lVert u_{a}\right\rVert_{r}\leq C,\quad\forall\,a\in(0,a_{1}).$ (3.7)
Proof.
Uniform bound of the positive part: For $M>0$, define $(u_{a}^{+})_{M}=\min\{(u_{a})^{+},M\}$ for every $a\in(0,a_{1})$. Clearly $(u_{a}^{+})_{M}\geq 0$ and $(u_{a}^{+})_{M}\in L^{\infty}({\mathbb{R}}^{N})\cap{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Fixed $\sigma\geq 1$, define $\phi=((u_{a}^{+})_{M})^{\sigma}$. Now we claim that $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. First, we recall the following inequality from [30, (2.4), Page 1359]: for any $\alpha,\beta\in\mathbb{R}$ and $\sigma\geq 1$, we have
$\left||\alpha|^{\sigma-1}\alpha-|\beta|^{\sigma-1}\beta\right|\leq\sigma\left(|\alpha|^{\sigma-1}+|\beta|^{\sigma-1}\right)|\alpha-\beta|.$ (A)
Taking $\alpha=(u_{a}^{+})_{M}(x)$ and $\beta=(u_{a}^{+})_{M}(y)$ into (A) and using the fact that $0\leq(u_{a}^{+})_{M}\in L^{\infty}({\mathbb{R}}^{N})$, we get
$\left|\phi(x)-\phi(y)\right|$ $\leq\sigma\left(|(u_{a}^{+})_{M}(x)|^{\sigma-1}+|(u_{a}^{+})_{M}(y)|^{\sigma-1}\right)\left|(u_{a}^{+})_{M}(x)-(u_{a}^{+})_{M}(y)\right|$
$\leq 2\sigma\|(u_{a}^{+})_{M}\|_{\infty}^{\sigma-1}\left|(u_{a}^{+})_{M}(x)-(u_{a}^{+})_{M}(y)\right|.$
Since $(u_{a}^{+})_{M}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we have $\left\lVert(u_{a}^{+})_{M}\right\rVert_{s,p}<\infty$ and which further implies $\left\lVert\phi\right\rVert_{s,p}<\infty$ i.e.,
$\left\lVert\phi\right\rVert_{s,p}^{p}=$ $\iint_{{\mathbb{R}}^{2N}}\frac{\lvert\phi(x)-\phi(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}\leq\left(2\sigma\|(u_{a}^{+})_{M}\|_{\infty}^{\sigma-1}\right)^{p}\iint_{{\mathbb{R}}^{2N}}\frac{\lvert(u_{a}^{+})_{M}(x)-(u_{a}^{+})_{M}(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}<\infty.$
Also, $(u_{a}^{+})_{M}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{\infty}({\mathbb{R}}^{N})$ implies $(u_{a}^{+})_{M}\in L^{p_{s}^{*}}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, which further implies $(u_{a}^{+})_{M}\in L^{r}({\mathbb{R}}^{N})$ for all $r\in[p_{s}^{*},\infty]$. Hence, we get $\phi=(u_{a}^{+})_{M}^{\sigma}\in L^{p_{s}^{*}}({\mathbb{R}}^{N})$. By definition of ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ given in (1.2), we conclude that $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$.
By Theorem 3.1, $u_{a}$ is a weak solution of (SP). Taking $\phi$ as a test function, we write
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u_{a}(x)-u_{a}(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{{\mathbb{R}}^{N}}g(x)f_{a}(u_{a})\phi(x)\,{\rm d}x.$
Using ($\bf i{2}$), Remark 3.2, and ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, we have the following lower bound of the left hand side:
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u_{a}(x)-u_{a}(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$~{}~{}\geq\frac{p^{p}}{(\sigma+p-1)^{p}}\iint_{{\mathbb{R}}^{2N}}\frac{\left|\big{(}(u_{a}^{+})_{M}(x)\big{)}^{\frac{\sigma+p-1}{p}}-\big{(}(u_{a}^{+})_{M}(y)\big{)}^{\frac{\sigma+p-1}{p}}\right|^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$~{}~{}~{}\geq\frac{C(N,s,p)p^{p}}{(\sigma+p-1)^{p}}\left(\int_{{\mathbb{R}}^{N}}\left(\big{(}(u_{a}^{+})_{M}(x)\big{)}^{\frac{\sigma+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}.$
Hence we get
$\frac{C(N,s,p)p^{p}}{(\sigma+p-1)^{p}}\left(\int_{{\mathbb{R}}^{N}}\left(\big{(}(u_{a}^{+})_{M}(x)\big{)}^{\frac{\sigma+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a}(u_{a})\rvert(u_{a}^{+}(x))^{\sigma}\,{\rm d}x.$
Taking the limit as $M\rightarrow\infty$ and using the monotone convergence theorem, we get
$\frac{C(N,s,p)p^{p}}{(\sigma+p-1)^{p}}\left(\int_{{\mathbb{R}}^{N}}\left((u_{a}^{+}(x))^{\frac{\sigma+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a}(u_{a})\rvert(u_{a}^{+}(x))^{\sigma}\,{\rm d}x.$ (3.8)
Step 1: In this step, for $\sigma=p^{*}_{s}$, we show that the set $\{(u_{a}^{+})^{\sigma+p-1}:a\in(0,a_{1})\}$ is uniformly bounded on $L^{\frac{\sigma}{p}}({\mathbb{R}}^{N})$. From ($\bf\tilde{f{1}}$) and ($\bf f{1}$), for any $\epsilon>0$, there exists $C=C({\epsilon},a_{1})$ such that $\lvert f_{a}(u_{a})\rvert\leq C+\epsilon\lvert u_{a}\rvert^{p^{*}_{s}-1}$ for all $a\in(0,a_{1})$. We can also write the previous inequality as follows:
$\lvert f_{a}(u_{a})\rvert\leq C+\epsilon(u_{a}^{+}+u_{a}^{-})^{p^{*}_{s}-1}=C+\epsilon\left((u_{a}^{+})^{p^{*}_{s}-1}+(u_{a}^{-})^{p^{*}_{s}-1}\right).$ (e0)
Since $u_{a}^{+}$ and $u_{a}^{-}$ have disjoint supports, we estimate the right hand side of (3.8) using (e0) as follows:
$\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a}(u_{a})\rvert(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x\leq C\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x+{\epsilon}\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{2p^{*}_{s}-1}\,{\rm d}x.$ (e1)
First we estimate the first integral on the right hand side of (e1). We use the continuous embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, the uniform boundedness of $\{u_{a}:a\in(0,a_{1})\}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ (Theorem 3.1), and Remark 3.2 to deduce that
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x\leq C(N,s,p)\left\lVert g\right\rVert_{\infty}\left\lVert u_{a}^{+}\right\rVert_{s,p}^{p^{*}_{s}}\leq C(N,s,p,f,g,a_{1}),\quad\forall\,a\in(0,a_{1}).$ (e2)
In order to estimate the second integral on the right hand side of (e1), we use the Hölder’s inequality with the conjugate pair $(\frac{p^{*}_{s}}{p^{*}_{s}-p},\frac{p^{*}_{s}}{p})$ and we get
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{2p^{*}_{s}-1}\,{\rm d}x$ $\leq\left\lVert g\right\rVert_{\infty}\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}-p}(u_{a}^{+}(x))^{p^{*}_{s}+p-1}\,{\rm d}x$
$\leq\left\lVert g\right\rVert_{\infty}\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p^{*}_{s}-p}{p^{*}_{s}}}\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}.$
Now ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$ and again the uniform boundedness of $\{u_{a}\}$ and Remark 3.2 yield
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{2p^{*}_{s}-1}\,{\rm d}x\leq C(N,s,p,f,g,a_{1})\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}.$ (e3)
Combining (3.8), (e1), (e2) and (e3), we obtain
$\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\left(\frac{p^{*}_{s}+p-1}{p}\right)^{p}\left(C_{1}+C_{2}{\epsilon}\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\right),$ (3.9)
where $C_{1}=C_{1}(N,s,p,{\epsilon},f,g,a_{1})$ and $C_{2}=C_{2}(N,s,p,f,g,a_{1})$. Now we choose ${\epsilon}>0$ such that
${\epsilon}\left(\frac{p^{*}_{s}+p-1}{p}\right)^{p}C_{2}<\frac{1}{2}.$
Using the above choice of $\epsilon>0$, the inequality (3.9) reduces to
$\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\left(\frac{p^{*}_{s}+p-1}{p}\right)^{p}C_{1}(N,s,p,f,g,a_{1}),\quad\forall\,a\in(0,a_{1}).$
Thus the set $\{(u_{a}^{+})^{p^{*}_{s}+p-1}:a\in(0,a_{1})\}$ is uniformly bounded on $L^{\frac{p^{*}_{s}}{p}}({\mathbb{R}}^{N})$.
Step 2: In this step, we consider $\sigma>p^{*}_{s}$. Using $\lvert f_{a}(t)\rvert\leq C(f,a_{1})(1+\lvert t\rvert^{p^{*}_{s}-1});t\in{\mathbb{R}}$ (by (2.3)) we have
$\lvert f_{a}(u_{a})\rvert\leq C(f,a_{1})\left(1+\left((u_{a}^{+})^{p^{*}_{s}-1}+(u_{a}^{-})^{p^{*}_{s}-1}\right)\right),$
and hence (3.8) yields
$\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\frac{(\sigma+p-1)^{p}}{C(N,s,p)p^{p}}C(f,a_{1})\int_{{\mathbb{R}}^{N}}g(x)\left(1+(u_{a}^{+}(x))^{p^{*}_{s}-1}\right)(u_{a}^{+}(x))^{\sigma}\,{\rm d}x.$ (3.10)
Set $m_{1}:=\frac{p^{*}_{s}(p^{*}_{s}-1)}{\sigma-1}$ and $m_{2}:=\sigma-m_{1}$. Notice that $m_{1}<p^{*}_{s}$. Applying Young’s inequality with the conjugate pair $(\frac{p^{*}_{s}}{m_{1}},\frac{p^{*}_{s}}{p^{*}_{s}-m_{1}})$ we get
$(u_{a}^{+}(x))^{\sigma}=(u_{a}^{+}(x))^{m_{1}}(u_{a}^{+}(x))^{m_{2}}\leq\frac{m_{1}}{p^{*}_{s}}(u_{a}^{+}(x))^{p^{*}_{s}}+\frac{p^{*}_{s}-m_{1}}{p^{*}_{s}}(u_{a}^{+}(x))^{\frac{p^{*}_{s}m_{2}}{p^{*}_{s}-m_{1}}},$ (3.11)
where we further observe that
$m_{1}=\frac{p^{*}_{s}(p^{*}_{s}-1)}{\sigma-1}\Longleftrightarrow\frac{p^{*}_{s}m_{2}}{p^{*}_{s}-m_{1}}=p^{*}_{s}+\sigma-1.$ (3.12)
Therefore, using ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$ and the uniform boundedness of $\{u_{a}\}$,
$\int_{{\mathbb{R}}^{N}}g(x)(u_{a}^{+}(x))^{\sigma}\,{\rm d}x$ $\leq\left\lVert g\right\rVert_{\infty}\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}}\,{\rm d}x+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}+\sigma-1}\,{\rm d}x\right)$
$\leq C(N,s,p,f,g,a_{1})\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}+\sigma-1}\,{\rm d}x\right).$
Hence for every $\sigma>p^{*}_{s}$, (3.10) yields
$\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\left(\frac{\sigma+p-1}{p}\right)^{p}C\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}+\sigma-1}\,{\rm d}x\right),$
where $C=C(N,s,p,f,g,a_{1})$. From the above inequality, we get
$\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}$ $\leq 1+\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}$
$\leq 1+C\left(\frac{\sigma+p-1}{p}\right)^{p}\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}+\sigma-1}\,{\rm d}x\right)$
$\leq\left(1+C\left(\sigma+p-1\right)^{p}\right)\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}+\sigma-1}\,{\rm d}x\right)$
$\leq\widetilde{C}\left(\sigma+p-1\right)^{p}\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{p^{*}_{s}+\sigma-1}\,{\rm d}x\right),$ (3.13)
where $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{1})=C+(p-1)^{-p}$. We consider the sequence $\{\sigma_{j}\}$ such that
$\sigma_{1}=p^{*}_{s},\sigma_{2}=1+\frac{p^{*}_{s}}{p}(\sigma_{1}-1),\cdots,\sigma_{j+1}=1+\frac{p^{*}_{s}}{p}(\sigma_{j}-1).$
Notice that $p^{*}_{s}+\sigma_{j+1}-1=\frac{\sigma_{j}+p-1}{p}p^{*}_{s}$ and $\sigma_{j+1}-1=(\frac{p^{*}_{s}}{p})^{j}(\sigma_{1}-1)$. Hence using (3), we have
$\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma_{j+1}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}(\sigma_{j+1}-1)}}\leq$ $\left(\widetilde{C}\left(\sigma_{j+1}+p-1\right)^{p}\right)^{\frac{1}{\sigma_{j+1}-1}}$
$\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma_{j}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}(\sigma_{j}-1)}}.$
Set $D_{j}:=\left(1+\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma_{j}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}(\sigma_{j}-1)}}$. Set $\eta_{j}=\sigma_{j}+p-1$. We iterate the above inequality to get
$D_{j+1}\leq\left({\left(\,\widetilde{C}\,\right)}^{\sum_{k=2}^{j+1}\frac{1}{\sigma_{k}-1}}\left(\prod_{k=2}^{j+1}\eta_{k}^{\frac{1}{\eta_{k}-p}}\right)^{p}\right)D_{1}.$
In view of Step 1, $D_{1}\leq C$ for some $C=C(N,s,p,f,g,a_{1})$. Moreover,
$D_{j+1}\geq\left(\left(\int_{{\mathbb{R}}^{N}}(u_{a}^{+}(x))^{\frac{\sigma_{j+1}+p-1}{p}p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{(\sigma_{j+1}+p-1)p^{*}_{s}}}\right)^{\frac{\sigma_{j+1}+p-1}{\sigma_{j+1}-1}}=\left\lVert u_{a}^{+}\right\rVert_{\frac{\sigma_{j+1}+p-1}{p}p^{*}_{s}}^{\frac{\eta_{j+1}}{\eta_{j+1}-p}}.$
Therefore,
$\left\lVert u_{a}^{+}\right\rVert_{\frac{\sigma_{j+1}+p-1}{p}p^{*}_{s}}^{\frac{\eta_{j+1}}{\eta_{j+1}-p}}\leq\left({\left(\,\widetilde{C}\,\right)}^{\sum_{k=2}^{j+1}\frac{1}{\eta_{k}-p}}\left(\prod_{k=2}^{j+1}\eta_{k}^{\frac{1}{\eta_{k}-p}}\right)^{p}\right)C,\quad\forall\,a\in(0,a_{1}),$ (3.14)
where $C,\widetilde{C}$ are independent of $a$. Since, $\sigma_{j}\rightarrow\infty$, as $j\rightarrow\infty$, by interpolation argument we have $u_{a}^{+}\in L^{r}({\mathbb{R}}^{N})$ for every $r\in[p^{*}_{s},\infty)$, and moreover from (3.14), $\left\lVert u_{a}^{+}\right\rVert_{r}\leq C$ for all $a\in(0,a_{1})$ and $C=C(r,N,s,p,f,g,a_{1})$. Further,
$\sum_{k=2}^{\infty}\frac{1}{\eta_{k}-p}=\frac{1}{(\eta_{1}-p)}\sum_{k=2}^{\infty}\left(\frac{p}{p_{s}^{*}}\right)^{k-1}=\frac{p}{(p_{s}^{*}-1)(p_{s}^{*}-p)}\text{ and }$
$\prod_{k=2}^{\infty}\eta_{k}^{\frac{1}{\eta_{k}-p}}=\exp\left(\sum_{k=2}^{\infty}\frac{\log(\eta_{k})}{\eta_{k}-p}\right)=\exp\left(\frac{p}{(p^{*}_{s}-p)^{2}}\log\left(p\left(\frac{p^{*}_{s}(p^{*}_{s}-p)}{p}\right)^{p^{*}_{s}}\right)\right).$
Also observe that $\frac{\eta_{j+1}}{\eta_{j+1}-p}\rightarrow 1$ as $j\rightarrow\infty$. Therefore, taking the limit as $j\rightarrow\infty$ in (3.14) gives $u_{a}^{+}\in L^{\infty}({\mathbb{R}}^{N})$ and $\left\lVert u_{a}^{+}\right\rVert_{\infty}\leq C(N,s,p,f,g,a_{1})$ for all $a\in(0,a_{1})$.
Uniform bound of the negative part: For $M>0$, define $(u_{a}^{-})_{M}=\min\{u_{a}^{-},M\}$ for every $a\in(0,a_{1})$. Clearly $(u_{a}^{-})_{M}\geq 0$ and $(u_{a}^{-})_{M}\in L^{\infty}({\mathbb{R}}^{N})\cap{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. For $\sigma\geq 1$, we take $\phi=-((u_{a}^{-})_{M})^{\sigma}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ as a test function to get
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u_{a}(x)-u_{a}(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{{\mathbb{R}}^{N}}g(x)f_{a}(u_{a})\phi(x)\,{\rm d}x.$
Using ($\bf i{4}$), Remark 3.2, and ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, we have
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u_{a}(x)-u_{a}(y)\big{)}\big{(}(u_{a}^{-})_{M}(y))^{\sigma}-((u_{a}^{-})_{M}(x))^{\sigma}\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$~{}~{}\geq\frac{p^{p}}{(\sigma+p-1)^{p}}\iint_{{\mathbb{R}}^{2N}}\frac{\left|((u_{a}^{-})_{M}(x))^{\frac{\sigma+p-1}{p}}-((u_{a}^{-})_{M}(y))^{\frac{\sigma+p-1}{p}}\right|^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$~{}~{}~{}\geq\frac{C(N,s,p)p^{p}}{(\sigma+p-1)^{p}}\left(\int_{{\mathbb{R}}^{N}}\left((u_{a}^{-})_{M}(x))^{\frac{\sigma+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}.$
Hence we get
$\frac{C(N,s,p)p^{p}}{(\sigma+p-1)^{p}}\left(\int_{{\mathbb{R}}^{N}}\left((u_{a}^{-})_{M}(x)^{\frac{\sigma+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a}(u_{a})\rvert(u_{a}^{-}(x))^{\sigma}\,{\rm d}x.$
Taking the limit as $M\rightarrow\infty$ and using the monotone convergence theorem, we get
$\frac{C(N,s,p)p^{p}}{(\sigma+p-1)^{p}}\left(\int_{{\mathbb{R}}^{N}}\left((u_{a}^{-}(x))^{\frac{\sigma+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a}(u_{a})\rvert(u_{a}^{-}(x))^{\sigma}\,{\rm d}x.$ (3.15)
Now, using (3.15) and following the same procedure as above, we can conclude that for every $r\in[p^{*}_{s},\infty]$, $u_{a}^{-}\in L^{r}({\mathbb{R}}^{N})$ and $\left\lVert u_{a}^{-}\right\rVert_{r}\leq C(r,N,s,p,f,g,a_{1})$ for all $a\in(0,a_{1})$.
Further, for $r\in[p^{*}_{s},\infty]$,
$\left\lVert u_{a}\right\rVert_{r}=\left\lVert u_{a}^{+}-u_{a}^{-}\right\rVert_{r}\leq\left\lVert u_{a}^{+}\right\rVert_{r}+\left\lVert u_{a}^{-}\right\rVert_{r}\leq C(r,N,s,p,f,g,a_{1}),\quad\forall\,a\in(0,a_{1}).$
Thus (3.7) holds. ∎
Now, we prove that the solution $\{u_{a}\}$ is uniformly bounded from below over several spaces.
Proposition 3.4.
Let $p\in(1,\frac{N}{s})$ and $f,g,a_{1}$ be given in Proposition 3.3. Then the following hold:
  1. (i)
    There exists $C_{1}>0$ such that $\|u_{a}\|_{s,p}\geq C_{1}$, for all $a\in(0,a_{1})$.
  2. (ii)
    There exist $a_{2}\in(0,a_{1})$ and $C_{2}>0$ such that $\|u_{a}\|_{\infty}\geq C_{2}$, for all $a\in(0,a_{2})$.
Proof.
(i) We notice that $F_{a}(t)\geq-a\lvert t\rvert$ for all $t\in{\mathbb{R}}$. For $\delta$ as given in Theorem 3.1, from (3.1) we write $I_{a}(u_{a})\geq\delta$, for all $a\in(0,a_{1})$. Using Proposition 2.2, we have
$\delta\leq I_{a}(u_{a})\leq\frac{1}{p}\|u_{a}\|_{s,p}^{p}+a\int_{{\mathbb{R}}^{N}}g(x)|u_{a}|\,{\rm d}x\leq\frac{1}{p}\|u_{a}\|_{s,p}^{p}+a_{1}C\|u_{a}\|_{s,p},$
where $C=C(N,s,p,g)$. Thus from the above inequality, there exists $C_{1}=C_{1}(N,s,p,g,a_{1},\delta)$ such that $\|u_{a}\|_{s,p}\geq C_{1}$, for all $a\in(0,a_{1})$.
(ii) For $\delta$ as given in Theorem 3.1, from (3.1) we write $I_{a}(u_{a})\geq\delta$ for all $a\in(0,a_{1})$. Further, we have
$\frac{\|u_{a}\|_{s,p}^{p}}{p}=I_{a}(u_{a})+\int_{{\mathbb{R}}^{N}}g(x)F_{a}(u_{a})\,{\rm d}x$ $\geq\delta-a\int_{{\mathbb{R}}^{N}}g(x)\lvert u_{a}\rvert\,{\rm d}x.$
For every $a\in(0,a_{1})$, using the continuous embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{1}({\mathbb{R}}^{N},g)$ (Proposition 2.2) with embedding constant $C_{2}=C_{2}(N,s,p,g)$ and the uniform boundedness of $\{u_{a}\}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ (Theorem 3.1), we obtain
$\frac{\|u_{a}\|_{s,p}^{p}}{p}\geq\delta-aC_{2}\|u_{a}\|_{s,p}\geq\delta-aC_{2}C=\delta-aC_{3},$
where $C_{3}=C_{2}C$. Now if we choose $a_{2}$ such that $0<a_{2}<\min\{\frac{\delta}{C_{3}},a_{1}\}$, then
$\frac{\|u_{a}\|_{s,p}^{p}}{p}\geq\delta_{0}:=\delta-a_{2}C_{3}>0,\quad\forall\,a\in(0,a_{2}).$
Therefore, using $|f_{a}(u_{a})|\leq C(f,a_{2})(1+|u_{a}|^{p_{s}^{*}-1})$ and that $u_{a}\in L^{\infty}({\mathbb{R}}^{N})$, we obtain the following estimates:
$\delta_{0}\leq\frac{\|u_{a}\|_{s,p}^{p}}{p}=\frac{1}{p}\int_{{\mathbb{R}}^{N}}g(x)|f_{a}(u_{a})u_{a}|\,{\rm d}x\leq C(f,a_{2})\|g\|_{1}(\|u_{a}\|_{\infty}+\|u_{a}\|_{\infty}^{p_{s}^{*}}).$ (3.16)
Thus we can conclude from (3.16) that there exists $C_{2}=C_{2}(N,s,p,f,g,a_{2},\delta)$ such that $\|u_{a}\|_{\infty}\geq C_{2}$ for all $a\in(0,a_{2})$. ∎
Proposition 3.5.
Let $p\in(1,\frac{N}{s})$ and $f,g$ be given in Proposition 3.3. Given a sequence $a_{n}\rightarrow 0$, there exists $\tilde{u}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ such that $u_{a_{n}}\rightarrow\tilde{u}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ as $n\rightarrow\infty$. Moreover, $\tilde{u}$ is a weak solution of
$(-\Delta)_{p}^{s}u=g(x)f_{0}(u)\text{ in }\mathbb{R}^{N},$ (3.17)
where $f_{0}(t)=f(t)$ for $t\geq 0$ and $f_{0}(t)=0$ for $t\leq 0$. Further, $\tilde{u}\in L^{r}({\mathbb{R}}^{N})$ for every $r\in[p^{*}_{s},\infty]$. Furthermore, $\tilde{u}\in\mathcal{C}({\mathbb{R}}^{N})$.
Proof.
Since $a_{n}\rightarrow 0$ as $n\rightarrow\infty$, there exists $n_{1}\in{\mathbb{N}}$ such that $a_{n}<a_{2}$ for all $n\geq n_{1}$. For each $n\geq n_{1}$, $u_{a_{n}}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ is a critical point of $I_{a_{n}}$. Moreover, by Theorem 3.1, the following hold up to a subsequence:
$I_{a_{n}}(u_{a_{n}})\rightarrow c\text{ in }{\mathbb{R}}\text{ as }n\rightarrow\infty,I^{\prime}_{a_{n}}(u_{a_{n}})=0\text{ and }\left\lVert u_{a_{n}}\right\rVert_{s,p}\leq C\quad\forall\,n\geq n_{1}.$
Therefore, $\{u_{a_{n}}\}$ is a bounded P-S sequence, and hence up to subsequence $u_{a_{n}}\rightarrow\tilde{u}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. This implies $u_{a_{n}}\rightarrow\tilde{u}$ in $L^{p^{*}_{s}}({\mathbb{R}}^{N})$ and up to a further subsequence $u_{a_{n}}(x)\rightarrow\tilde{u}(x)$ a.e. in ${\mathbb{R}}^{N}$. Now for every $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ with $\phi\geq 0$ we show that
$\int_{{\mathbb{R}}^{N}}g(x)f_{a_{n}}(u_{a_{n}})\phi(x)\,{\rm d}x\rightarrow\int_{{\mathbb{R}}^{N}}g(x)f_{0}(\tilde{u})\phi(x)\,{\rm d}x,\text{as }n\rightarrow\infty.$
We split
$\lvert f_{a_{n}}(u_{a_{n}})-f_{0}(\tilde{u})\rvert\leq\lvert f_{a_{n}}(u_{a_{n}})-f_{0}(u_{a_{n}})\rvert+\lvert f_{0}(u_{a_{n}})-f_{0}(\tilde{u})\rvert.$
Using the continuity of $K_{0}^{\prime}$, $\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{0}(u_{a_{n}})-f_{0}(\tilde{u})\rvert\phi(x)\rightarrow 0$, as $n\rightarrow\infty$. Further, $\lvert f_{a_{n}}(u_{a_{n}})-f_{0}(u_{a_{n}})\rvert\leq a_{n}$. Therefore,
$\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{a_{n}}(u_{a_{n}})-f_{0}(\tilde{u})\rvert\phi(x)\,{\rm d}x$
$\leq a_{n}\int_{{\mathbb{R}}^{N}}g(x)\phi(x)\,{\rm d}x+\int_{{\mathbb{R}}^{N}}g(x)\lvert f_{0}(u_{a_{n}})-f_{0}(\tilde{u})\rvert\phi(x)\,{\rm d}x\rightarrow 0,\text{ as }n\rightarrow\infty.$
From the weak formulation
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u_{a_{n}}(x)-u_{a_{n}}(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{{\mathbb{R}}^{N}}g(x)f_{a_{n}}(u_{a_{n}})\phi(x)\,{\rm d}x.$
Taking the limit as $n\rightarrow\infty$ gives
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}\tilde{u}(x)-\tilde{u}(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{{\mathbb{R}}^{N}}g(x)f_{0}(\tilde{u})\phi(x)\,{\rm d}x,~{}~{}\forall\,\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})},\phi\geq 0.$
Now, for any $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we write $\phi=\phi^{+}-\phi^{-}$ where we see that the above identity holds for both $\phi^{+}$ and $\phi^{-}$. Thus
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}\tilde{u}(x)-\tilde{u}(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{{\mathbb{R}}^{N}}g(x)f_{0}(\tilde{u})\phi(x)\,{\rm d}x,~{}~{}\forall\,\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})},$ (3.18)
which implies that $\tilde{u}$ is a weak solution to (3.17). Now using a similar set of arguments as in the proof of Proposition 3.3, we obtain $\tilde{u}\in L^{r}({\mathbb{R}}^{N})$ for every $r\in[p^{*}_{s},\infty]$. Next, we show that $\tilde{u}\in\mathcal{C}({\mathbb{R}}^{N})$. Let $\Omega\subset{\mathbb{R}}^{N}$ be an open and bounded Lipschitz set. Let $\phi\in W^{s,p}(\Omega)$ and $\phi$ is compactly supported in $\Omega$. Define $\tilde{\phi}(x)=\phi(x)$ for $x\in\Omega$, and $\tilde{\phi}(x)=0$ for $x\in{\mathbb{R}}^{N}\setminus\Omega$. According to [27, Theorem 5.1], $\tilde{\phi}\in W^{s,p}({\mathbb{R}}^{N})$. Using $W^{s,p}({\mathbb{R}}^{N})\hookrightarrow L^{p_{s}^{*}}({\mathbb{R}}^{N})$[27, Theorem 6.5], it follows that $\tilde{\phi}\in L^{p_{s}^{*}}({\mathbb{R}}^{N})$. Therefore, based on the characterization of ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ in (1.2), we conclude that $\tilde{\phi}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. Now $\tilde{u}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{\infty}({\mathbb{R}}^{N})$. Using the embeddings ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$ and $L^{p^{*}_{s}}(\Omega)\hookrightarrow L^{p}(\Omega)$, observe that ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\subset W^{s,p}_{\text{loc}}(\Omega)$. Also from the definition, $L^{\infty}({\mathbb{R}}^{N})\subset L^{p-1}_{sp}({\mathbb{R}}^{N})$. Therefore, in view of (3.18) we see that $\tilde{u}\in W^{s,p}_{\text{loc}}(\Omega)\cap L^{p-1}_{sp}({\mathbb{R}}^{N})$ satisfies the following identity:
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}\tilde{u}(x)-\tilde{u}(y)\big{)}\big{(}\tilde{\phi}(x)-\tilde{\phi}(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$ $=\int_{{\mathbb{R}}^{N}}g(x)f_{0}(\tilde{u})\tilde{\phi}(x)\,{\rm d}x,$
$=\int_{\Omega}g(x)f_{0}(\tilde{u})\phi(x)\,{\rm d}x,~{}~{}\forall\phi\in W^{s,p}(\Omega),\text{supp}(\phi)\subset\Omega.$
Moreover, $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, $\lvert f_{0}(\tilde{u})\rvert\leq\epsilon\lvert\tilde{u}\rvert^{p-1}+C(f,\epsilon)\lvert\tilde{u}\rvert^{\gamma-1}$ (by (2.2)), and $\tilde{u}\in L^{\infty}({\mathbb{R}}^{N})$. Hence, $g(x)f_{0}(\tilde{u})\in L^{q}({\mathbb{R}}^{N})$ for $q>\frac{N}{sp}$. Thus, applying [10, Theorem 1.4] for $p\geq 2$ and [29, Theorem 1.2] for $21<p<$ 2 over $\tilde{u}$, where
$(-\Delta)_{p}^{s}\tilde{u}=g(x)f_{0}(\tilde{u})\text{ in }\Omega,$
we conclude that $\tilde{u}\in\mathcal{C}_{\text{loc}}^{\delta}(\Omega)$ for some $\delta\in(0,1)$. In particular, $\tilde{u}\in\mathcal{C}_{\text{loc}}(\Omega)$ and hence we can get $\tilde{u}\in\mathcal{C}(\Omega)$ for every bounded open Lipschitz set $\Omega\subset{\mathbb{R}}^{N}$. Next, for any compact set $K\subset{\mathbb{R}}^{N}$, we have $K\subset\Omega$ for some bounded open Lipschitz set $\Omega$. Thus, $\tilde{u}\in\mathcal{C}(K)$ as well which implies $\tilde{u}\in\mathcal{C}_{\text{loc}}({\mathbb{R}}^{N})$. Therefore, $\tilde{u}\in\mathcal{C}({\mathbb{R}}^{N})$. ∎
Next, for a sequence $\{a_{n}\}$ going to zero, we prove the uniform convergence of $\{u_{a_{n}}\}$ over ${\mathbb{R}}^{N}$.
Proposition 3.6.
Let $p\in(\frac{2N}{N+2s},\frac{N}{s})$. Consider $a_{n},\tilde{u}$ given in Proposition 3.5. Then
$\left\lVert u_{a_{n}}-\tilde{u}\right\rVert_{\infty}\rightarrow 0,\text{ as }n\rightarrow\infty.$
Proof.
Using $a_{n}\rightarrow 0$, there exists $n_{1}\in{\mathbb{N}}$ such that $a_{n}<a_{2}$ for all $n\geq n_{1}$. Now for $n\geq n_{1}$, $I_{a_{n}}^{\prime}(u_{a_{n}})=0$, and hence the following identity holds:
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi(u_{a_{n}}(x)-u_{a_{n}}(y))(\phi(x)-\phi(y))}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{{\mathbb{R}}^{N}}g(x)f_{{a_{n}}}(u_{a_{n}})\phi(x)\,{\rm d}x,\quad\forall\,\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}.$
For brevity we denote $u_{n}:=u_{a_{n}}$ and $f_{n}(\cdot):=f_{a_{n}}(\cdot)$. Applying Proposition 3.5, $u_{n}\rightarrow\tilde{u}$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ and
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi(\tilde{u}(x)-\tilde{u}(y))(\phi(x)-\phi(y))}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{{\mathbb{R}}^{N}}g(x)f_{0}(\tilde{u})\phi(x)\,{\rm d}x,\quad\forall\,\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}.$
Subtracting the above identities, we get
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi(u_{n}(x)-u_{n}(y))-\Phi(\tilde{u}(x)-\tilde{u}(y))}{|x-y|^{N+sp}}(\phi(x)-\phi(y))\,{\mathrm{d}x\mathrm{d}y}$
$=\int_{{\mathbb{R}}^{N}}g(x)\left(f_{n}(u_{n})-f_{0}(\tilde{u})\right)\phi(x)\,{\rm d}x.$ (3.19)
Now we define $w_{n}=u_{n}-\tilde{u}$. Since $u_{n},\tilde{u}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{r}({\mathbb{R}}^{N})$ (Proposition 3.3 and Proposition 3.5), we have $w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\cap L^{r}({\mathbb{R}}^{N})$ for every $r\in[p^{*}_{s},\infty]$. Taking $\alpha=w_{n}(x)$ and $\beta=w_{n}(y)$ into the inequality (A) and following the same arguments in the proof of Proposition 3.3, where it is shown that $\phi=((u_{a}^{+})_{M})^{\sigma}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$, we get $\lvert w_{n}\rvert^{q-1}w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ for every $q\geq 1$.
For $p\geq 2$: We use [31, Lemma 2.3] to obtain the following lower bound of the right hand side of (3):
$\left\lVert\lvert w_{n}\rvert^{\frac{q+p-1}{p}-1}w_{n}\right\rVert_{s,p}^{p}\leq C(p)q^{p-1}$ $\iint_{{\mathbb{R}}^{2N}}\frac{\Phi(u_{n}(x)-u_{n}(y))-\Phi(\tilde{u}(x)-\tilde{u}(y))}{|x-y|^{N+sp}}$
$\left(\lvert w_{n}(x)\rvert^{q-1}w_{n}(x)-(\lvert w_{n}(y)\rvert^{q-1}w_{n}(y)\right)\,{\mathrm{d}x\mathrm{d}y}.$
Hence using the embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$ and choosing $\phi=\lvert w_{n}\rvert^{q-1}w_{n}$ in (3) we obtain
$C(N,s,p)\left(\int_{{\mathbb{R}}^{N}}\left(\lvert w_{n}(x)\rvert^{\frac{q+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq\left\lVert\lvert w_{n}\rvert^{\frac{q+p-1}{p}-1}w_{n}\right\rVert_{s,p}^{p}$
$\leq C(p)q^{p-1}\int_{{\mathbb{R}}^{N}}g(x)\left|f_{n}(u_{n})-f_{0}(\tilde{u})\right||w_{n}(x)|^{q}\,{\rm d}x.$ (3.20)
Define $g_{n}(t):=f_{n}(t+\tilde{u})-f_{0}(\tilde{u}).$ Using (2.3), we have $\lvert f_{n}(t)\rvert\leq C(f,a_{2})(1+\lvert t\rvert^{p^{*}_{s}-1}),\lvert f_{0}(t)\rvert\leq C(f)(1+\lvert t\rvert^{p^{*}_{s}-1});t\in{\mathbb{R}}$. By noting that $g_{n}(w_{n})=f_{n}(u_{n})-f_{0}(\tilde{u})$,
$\lvert g_{n}(w_{n})\rvert\leq\lvert f_{n}(u_{n})\rvert+\lvert f_{0}(\tilde{u})\rvert\leq C(f,a_{2})\left(1+\lvert u_{n}\rvert^{p^{*}_{s}-1}+\lvert\tilde{u}\rvert^{p^{*}_{s}-1}\right)\leq C\left(1+\lvert w_{n}\rvert^{p^{*}_{s}-1}\right),$ (3.21)
where $C$ does not depend on $a_{n}$. Therefore, (3) yields
$\left(\int_{{\mathbb{R}}^{N}}\left(\lvert w_{n}(x)\rvert^{\frac{q+p-1}{p}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}$ $\leq C(N,s,p,f,a_{2})q^{p-1}$
$\int_{{\mathbb{R}}^{N}}g(x)\left(1+\lvert w_{n}(x)\rvert^{p^{*}_{s}-1}\right)\lvert w_{n}(x)\rvert^{q}\,{\rm d}x.$ (3.22)
Let us consider
$\bar{q}>\frac{N(p_{s}^{*}-1)}{sp}\;\text{ and }\;\bar{r}=\frac{\bar{q}}{p^{*}_{s}-1}.$ (3.23)
Observe that $\frac{N(p_{s}^{*}-1)}{sp}>p_{s}^{*}$. Hence $\{w_{n}\}$ is bounded in $L^{\bar{q}}({\mathbb{R}}^{N})$ (by Proposition 3.3 and Proposition 3.5). For the conjugate pair $(\bar{r},\bar{r}^{\prime})$, define
$G_{n}(x,w_{n}):=(g(x))^{\frac{1}{\bar{r}}}\left(1+\lvert w_{n}(x)\rvert^{p^{*}_{s}-1}\right)\,\text{ and }\,H_{n}(x,w_{n}):=(g(x))^{\frac{1}{\bar{r}^{\prime}}}\lvert w_{n}(x)\rvert^{q}.$ (3.24)
Since $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, we have
$\int_{{\mathbb{R}}^{N}}\lvert G_{n}(x,w_{n})\rvert^{\bar{r}}\,{\rm d}x\leq 2^{\bar{r}-1}\int_{{\mathbb{R}}^{N}}g(x)\,{\rm d}x+2^{\bar{r}-1}\left\lVert g\right\rVert_{\infty}\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{\bar{q}}\,{\rm d}x\leq C(N,s,p,f,g,a_{2}).$
If $q\bar{r}^{\prime}<p_{s}^{*}$, then applying Proposition 2.2, $\int_{{\mathbb{R}}^{N}}\lvert H_{n}(x,w_{n})\rvert^{\bar{r}^{\prime}}\,{\rm d}x=\int_{{\mathbb{R}}^{N}}g(x)\lvert w_{n}(x)\rvert^{q\bar{r}^{\prime}}\,{\rm d}x<\infty$. If $q\bar{r}^{\prime}\geq p_{s}^{*}$, then using boundedness of $\{w_{n}\}$ in $L^{q\bar{r}^{\prime}}({\mathbb{R}}^{N})$ (by Proposition 3.3 and Proposition 3.5), $\int_{{\mathbb{R}}^{N}}\lvert H_{n}(x,w_{n})\rvert^{\bar{r}^{\prime}}\,{\rm d}x\leq\left\lVert g\right\rVert_{\infty}\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{q\bar{r}^{\prime}}\,{\rm d}x<\infty.$ Hence we apply the Hölder’s inequality with the conjugate pair $(\bar{r},\bar{r}^{\prime})$ to get from (3) that
$\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{\frac{(q+p-1)p^{*}_{s}}{p}}\,{\rm d}x\right)^{\frac{p}{p^{*}_{s}}}\leq Cq^{p-1}\int_{{\mathbb{R}}^{N}}G_{n}(x,w_{n})H_{n}(x,w_{n})\,{\rm d}x$
$\leq Cq^{p-1}\left(\int_{{\mathbb{R}}^{N}}\lvert G_{n}(x,w_{n})\rvert^{\bar{r}}\,{\rm d}x\right)^{\frac{1}{\bar{r}}}\left(\int_{{\mathbb{R}}^{N}}\lvert H_{n}(x,w_{n})\rvert^{\bar{r}^{\prime}}\,{\rm d}x\right)^{\frac{1}{\bar{r}^{\prime}}}$
$\leq\widetilde{C}q^{p-1}\left(\int_{{\mathbb{R}}^{N}}g(x)\lvert w_{n}(x)\rvert^{q\bar{r}^{\prime}}\,{\rm d}x\right)^{\frac{1}{\bar{r}^{\prime}}},$ (3.25)
where $C=C(N,s,p,f,a_{2})$ and $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. Define $\theta:=\frac{p^{*}_{s}}{p\bar{r}^{\prime}}.$ Using the definition of $\bar{r}$ in (3.23) it follows that $\theta>1$. We consider two sequences $\{l_{j}\}$ and $\{m_{j}\}$ by the following recursive process:
$l_{0}=p_{s}^{*},\quad l_{j+1}=\theta l_{j}+\frac{p^{*}_{s}(p-1)}{p},\quad m_{j}=\frac{l_{j}}{\bar{r}^{\prime}}.$
Observe that $l_{j}\geq p^{*}_{s}$ and $l_{j+1}=\frac{(m_{j}+p-1)p^{*}_{s}}{p}$ for each $j\in\mathbb{N}$, and $l_{j},m_{j}\rightarrow\infty$ as $j\rightarrow\infty$. Thus, for $q=m_{j}$ in (3) we obtain the following for all $n,j\in\mathbb{N}$:
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j+1}}\,{\rm d}x$ $\leq\left(\widetilde{C}m_{j}^{p-1}\right)^{\frac{p^{*}_{s}}{p}}\left(\int_{{\mathbb{R}}^{N}}g(x)\lvert w_{n}(x)\rvert^{m_{j}\bar{r}^{\prime}}\,{\rm d}x\right)^{\frac{p^{*}_{s}}{p\bar{r}^{\prime}}}$
$\leq\left(\widetilde{C}m_{j}^{p}\right)^{\theta\bar{r}^{\prime}}\left\lVert g\right\rVert_{\infty}^{\theta}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta}$
$\leq\left(\widetilde{C}m_{j}^{p}\right)^{\theta\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta},$ (3.26)
for some $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. By Lemma A.2 we observe that $m_{j}\sim\frac{\theta^{j}}{\bar{r}^{\prime}}$ as $j\rightarrow\infty$. Thus iterating on (3) we obtain
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x$ $\leq\prod_{i=0}^{j-1}\left(\widetilde{C}m_{i}^{p}\right)^{\theta^{j-i}\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{0}}\,{\rm d}x\right)^{\theta^{j}}$
$\leq\left(\widetilde{C}\right)^{\sum\limits_{i=0}^{j-1}\theta^{j-i}\bar{r}^{\prime}}\theta^{p\bar{r}^{\prime}\sum\limits_{i=0}^{j-1}i\theta^{j-i}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\theta^{j}}.$ (3.27)
Notice that $S_{1}:=\sum\limits_{i=0}^{\infty}\theta^{-i}<\infty$ and $S_{2}:=\sum\limits_{i=0}^{\infty}i\theta^{-i}<\infty$. Take $C>\max\{1,\widetilde{C}\}$. Then for all $n\geq n_{1},j\in{\mathbb{N}}$ using (3) we deduce
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\leq C^{\theta^{j}S_{1}}\theta^{p\bar{r}^{\prime}\theta^{j}S_{2}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\theta^{j}}.$
Since $w_{n}\rightarrow 0$ in $L^{p^{*}_{s}}({\mathbb{R}}^{N})$, there exists $n_{2}\in{\mathbb{N}}$ such that $\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}<1$ for all $n\geq n_{2}$. Therefore, for $n\geq\max\{n_{1},n_{2}\}$ from the above inequality, we obtain
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C^{\frac{\theta^{j}}{l_{j}}S_{1}}\theta^{p\bar{r}^{\prime}\frac{\theta^{j}}{l_{j}}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\frac{\theta^{j}}{l_{j}}}\leq C^{\beta_{2}S_{1}}\theta^{p\bar{r}^{\prime}\beta_{2}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\beta_{1}},$
where $\beta_{1},\beta_{2}>0$ have been chosen such that for all $j\in{\mathbb{N}}$, $\beta_{1}<\frac{\theta^{j}}{l_{j}}<\beta_{2}$ (see Lemma A.2). Therefore, there exists $C>1$ such that for all $n\geq\max\{n_{1},n_{2}\}$ and $j\in{\mathbb{N}}$ large enough
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{\beta_{1}p^{*}_{s}}$ (3.28)
Therefore, taking the limit as $j\rightarrow\infty$ in (3.28) and recalling that $w_{n}\in L^{\infty}({\mathbb{R}}^{N})$, we obtain for each $n\geq\max\{n_{1},n_{2}\}$,
$\left\lVert w_{n}\right\rVert_{\infty}\leq C\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{\beta_{1}p^{*}_{s}},$
where $C=C(N,s,p,f,g,a_{2})$. Finally, we take the limit as $n\rightarrow\infty$ in the above inequality and use $w_{n}\rightarrow 0$ in $L^{p^{*}_{s}}({\mathbb{R}}^{N})$ to get the required convergence.
For $\frac{2N}{N+2s}<p<2$: For any $q\geq 1$, we take $\lvert w_{n}\rvert^{q-1}w_{n}\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ as a test function in (3) and use [31, Lemma 2.4] to obtain
$\frac{\left\lVert\lvert w_{n}\rvert^{\frac{q+1}{2}-1}w_{n}\right\rVert_{s,p}^{2}}{\left(\left\lVert u_{n}\right\rVert_{s,p}^{p}+\left\lVert\tilde{u}\right\rVert_{s,p}^{p}\right)^{2-p}}$ $\leq Cq\iint_{{\mathbb{R}}^{2N}}\frac{\Phi(u_{n}(x)-u_{n}(y))-\Phi(\tilde{u}(x)-\tilde{u}(y))}{|x-y|^{N+sp}}$
$\quad\quad\quad\left(\lvert w_{n}(x)\rvert^{q-1}w_{n}(x)-(\lvert w_{n}(y)\rvert^{q-1}w_{n}(y)\right)\,{\mathrm{d}x\mathrm{d}y}$
$\leq Cq\int_{{\mathbb{R}}^{N}}g(x)\left|f_{n}(u_{n})-f_{0}(\tilde{u})\right||w_{n}(x)|^{q}\,{\rm d}x.$
Hence using ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, the uniform boundedness of $\{u_{n}\}$ over ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ (Theorem 3.1), and (3.21) we get
$\left(\int_{{\mathbb{R}}^{N}}\left(\lvert w_{n}(x)\rvert^{\frac{q+1}{2}}\right)^{p^{*}_{s}}\,{\rm d}x\right)^{\frac{2}{p^{*}_{s}}}\leq qC(N,s,p)\left(\left\lVert u_{n}\right\rVert_{s,p}^{p}+\left\lVert\tilde{u}\right\rVert_{s,p}^{p}\right)^{2-p}$
$\quad\quad\quad\int_{{\mathbb{R}}^{N}}g(x)\left|f_{n}(u_{n})-f_{0}(\tilde{u})\right||w_{n}(x)|^{q}\,{\rm d}x$
$\leq qC(N,s,p,f,g,a_{2})\int_{{\mathbb{R}}^{N}}g(x)\left(1+\lvert w_{n}(x)\rvert^{p^{*}_{s}-1}\right)\lvert w_{n}(x)\rvert^{q}\,{\rm d}x.$ (3.29)
Observe that $p>\frac{2N}{N+2s}\iff p_{s}^{*}>2$. For $\bar{q}$ as given in (3.23), we set
$\bar{r}=\frac{\bar{q}}{p^{*}_{s}-2}.$ (3.30)
For this $\bar{r}$ and its conjugate exponent $\bar{r}^{\prime}$, we consider $G_{n}$ and $H_{n}$ as defined in (3.24). Using the fact that $\bar{r}(p_{s}^{*}-1)>\bar{q}>p_{s}^{*}$ and boundedness of $\{w_{n}\}$ in $L^{\bar{r}(p_{s}^{*}-1)}({\mathbb{R}}^{N})$,
$\int_{{\mathbb{R}}^{N}}\lvert G_{n}(x,w_{n})\rvert^{\bar{r}}\,{\rm d}x\leq 2^{\bar{r}-1}\int_{{\mathbb{R}}^{N}}g(x)\,{\rm d}x+2^{\bar{r}-1}\left\lVert g\right\rVert_{\infty}\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{\bar{r}(p_{s}^{*}-1)}\,{\rm d}x\leq C(N,s,p,f,g,a_{2}).$
Moreover, following similar arguments as given earlier, $\int_{{\mathbb{R}}^{N}}\lvert H_{n}(x,w_{n})\rvert^{\bar{r}^{\prime}}\,{\rm d}x<\infty$. Thus applying the Hölder’s inequality with the conjugate pair $(\bar{r},\bar{r}^{\prime})$ we get from (3) that
$\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{\frac{(q+1)p^{*}_{s}}{2}}\,{\rm d}x\right)^{\frac{2}{p^{*}_{s}}}\leq Cq\int_{{\mathbb{R}}^{N}}G_{n}(x,w_{n})H_{n}(x,w_{n})\,{\rm d}x$
$\leq Cq\left(\int_{{\mathbb{R}}^{N}}\lvert G_{n}(x,w_{n})\rvert^{\bar{r}}\,{\rm d}x\right)^{\frac{1}{\bar{r}}}\left(\int_{{\mathbb{R}}^{N}}g(x)\lvert w_{n}(x)\rvert^{q\bar{r}^{\prime}}\,{\rm d}x\right)^{\frac{1}{\bar{r}^{\prime}}}$
$\leq\widetilde{C}q\left(\int_{{\mathbb{R}}^{N}}g(x)\lvert w_{n}(x)\rvert^{q\bar{r}^{\prime}}\,{\rm d}x\right)^{\frac{1}{\bar{r}^{\prime}}},$ (3.31)
where $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. Define $\theta:=\frac{p^{*}_{s}}{2\bar{r}^{\prime}}.$ Using the definition of $\bar{r}$ in (3.30) we can verify that $\theta>1$. We consider two sequences $\{l_{j}\}$ and $\{m_{j}\}$ by the following recursive process:
$l_{0}=p_{s}^{*},\quad l_{j+1}=\theta l_{j}+\frac{p^{*}_{s}}{2},\quad m_{j}=\frac{l_{j}}{\bar{r}^{\prime}}.$
Observe that $l_{j+1}=\frac{(m_{j}+1)p^{*}_{s}}{2}$ and $l_{j},m_{j}\rightarrow\infty$ as $j\rightarrow\infty$. Thus, for $q=m_{j}$ in (3) we obtain the following for all $n,j\in\mathbb{N}$:
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j+1}}\,{\rm d}x$ $\leq\left(\widetilde{C}m_{j}\right)^{\frac{p^{*}_{s}}{2}}\left(\int_{{\mathbb{R}}^{N}}g(x)\lvert w_{n}(x)\rvert^{m_{j}\bar{r}^{\prime}}\,{\rm d}x\right)^{\frac{p^{*}_{s}}{2\bar{r}^{\prime}}}$
$\leq\left(\widetilde{C}m_{j}^{2}\right)^{\theta\bar{r}^{\prime}}\left\lVert g\right\rVert_{\infty}^{\theta}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta}$
$\leq\left(\widetilde{C}m_{j}^{2}\right)^{\theta\bar{r}^{\prime}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\right)^{\theta},$ (3.32)
for some $\widetilde{C}=\widetilde{C}(N,s,p,f,g,a_{2})$. Iterating (3) and following a similar calculation, we can deduce that for all $n\geq n_{1},j\in{\mathbb{N}}$
$\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{l_{j}}\,{\rm d}x\leq C^{\theta^{j}S_{1}}\theta^{2\bar{r}^{\prime}\theta^{j}S_{2}}\left(\int_{{\mathbb{R}}^{N}}\lvert w_{n}(x)\rvert^{p^{*}_{s}}\,{\rm d}x\right)^{\theta^{j}},$
where $C>\max\{1,\widetilde{C}\}$. Since $w_{n}\rightarrow 0$ in $L^{p^{*}_{s}}({\mathbb{R}}^{N})$, there exists $n_{3}\in{\mathbb{N}}$ such that $\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}<1$ for all $n\geq n_{3}$. Therefore, for $n\geq\max\{n_{1},n_{3}\}$ we obtain
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C^{\frac{\theta^{j}}{l_{j}}S_{1}}\theta^{2\bar{r}^{\prime}\frac{\theta^{j}}{l_{j}}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\frac{\theta^{j}}{l_{j}}}\leq C^{\alpha_{2}S_{1}}\theta^{2\bar{r}^{\prime}\alpha_{2}S_{2}}\left(\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{p^{*}_{s}}\right)^{\alpha_{1}},$
where $\alpha_{1},\alpha_{2}$ have been chosen such that for all $j\in{\mathbb{N}}$, $\alpha_{1}<\frac{\theta^{j}}{l_{j}}<\alpha_{2}$ (see Lemma A.2). Finally, we find a $C>1$ such that for all $n\geq\max\{n_{1},n_{3}\}$ and $j\in{\mathbb{N}}$ large enough
$\left\lVert w_{n}\right\rVert_{l_{j}}\leq C\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{\alpha_{1}p^{*}_{s}}.$ (3.33)
Therefore, taking the limit as $j\rightarrow\infty$ in (3.33) we obtain for each $n\geq\max\{n_{1},n_{3}\}$,
$\left\lVert w_{n}\right\rVert_{\infty}\leq C\left\lVert w_{n}\right\rVert_{p^{*}_{s}}^{\alpha_{1}p^{*}_{s}},$
where $C=C(N,s,p,f,g,a_{2})$. Taking the limit as $n\rightarrow\infty$ in the above inequality, we get the required convergence. ∎
In the following proposition, we state a strong maximum principle for a nonlocal equation defined on ${\mathbb{R}}^{N}$. Our proof follows using the similar arguments given in [7, Theorem A.1]. For the sake of completeness, we sketch the proof.
Proposition 3.7 (Strong Maximum Principle).
Let $p\in(1,\infty)$ and $s\in(0,1)$. Let $f,g$ be given in Theorem 3.1. Assume that $u\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ is a weak supersolution of the following equation:
$(-\Delta)_{p}^{s}u=g(x)f(u)\text{ in }\mathbb{R}^{N},$
and $u\geq 0$ a.e. in ${\mathbb{R}}^{N}$. Then either $u\equiv 0$ or $u>0$ a.e. in ${\mathbb{R}}^{N}$.
Proof.
Take $\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ and $\phi\geq 0$. By the hypothesis, we have
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u(x)-u(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}\geq\int_{{\mathbb{R}}^{N}}g(x)f(u)\phi(x)\,{\rm d}x\geq 0.$
Let $K\subset\subset{\mathbb{R}}^{N}$ be any compact connected set. We first show that either $u\equiv 0$ or $u>0$ a.e. in $K$. Since $K$ is compact, we choose $x_{1},x_{2},\dots,x_{k}$ in ${\mathbb{R}}^{N}$ such that $K\subset\cup_{i=1}^{k}B_{\frac{r}{2}}(x_{i})$, and $\lvert B_{\frac{r}{2}}(x_{i})\cap B_{\frac{r}{2}}(x_{i+1})\rvert>0$ for each $i$. Suppose $u\equiv 0$ on a subset of $K$ with a positive measure. Then there exists $j\in\{1,\dots,k\}$ such that $\mathcal{A}=\{x\in B_{\frac{r}{2}}(x_{j}):u(x)=0\}$ has a positive measure, i.e., $|\mathcal{A}|>0$. We define
$F_{\delta}(x)=\log\left(1+\frac{u(x)}{\delta}\right)\text{ for }x\in B_{\frac{r}{2}}(x_{j})\text{ and }\delta>0.$
Clearly, $F_{\delta}\equiv 0$ on $\mathcal{A}$. Take $x\in B_{\frac{r}{2}}(x_{j})$ and $y\in\mathcal{A}$ with $y\neq x$. Then
$\lvert F_{\delta}(x)\rvert^{p}=\frac{\lvert F_{\delta}(x)-F_{\delta}(y)\rvert^{p}}{\lvert x-y\rvert^{N+sp}}\lvert x-y\rvert^{N+sp}.$
Integrating the above identity over $\mathcal{A}$ we get
$\lvert\mathcal{A}\rvert\lvert F_{\delta}(x)\rvert^{p}\leq\max_{x,y\in B_{\frac{r}{2}}(x_{j})}{\lvert x-y\rvert^{N+sp}}\int_{B_{\frac{r}{2}}(x_{j})}\left|\log\left(\frac{u(x)+\delta}{u(y)+\delta}\right)\right|^{p}\,\frac{{\mathrm{d}y}}{\lvert x-y\rvert^{N+sp}}.$
Further, the integration over $B_{\frac{r}{2}}(x_{j})$ yields
$\int_{B_{\frac{r}{2}}(x_{j})}\lvert F_{\delta}(x)\rvert^{p}\,{\rm d}x\leq\frac{r^{N+sp}}{\lvert\mathcal{A}\rvert}\iint_{B_{\frac{r}{2}}(x_{j})\times B_{\frac{r}{2}}(x_{j})}\left|\log\left(\frac{u(x)+\delta}{u(y)+\delta}\right)\right|^{p}\,\frac{{\mathrm{d}y}{\rm d}x}{\lvert x-y\rvert^{N+sp}}.$ (3.34)
Now on $B_{\frac{r}{2}}(x_{j})$ we use the Logarithmic energy estimate on $u$ (see [26, Lemma 1.3]), to get
$\iint_{B_{\frac{r}{2}}(x_{j})\times B_{\frac{r}{2}}(x_{j})}\left|\log\left(\frac{u(x)+\delta}{u(y)+\delta}\right)\right|^{p}\,\frac{{\mathrm{d}y}{\rm d}x}{\lvert x-y\rvert^{N+sp}}\leq C(N,s,p)r^{N-sp}.$ (3.35)
From (3.34) and (3.35) we get for every $\delta>0$ that
$\int_{B_{\frac{r}{2}}(x_{j})}\left|\log\left(1+\frac{u(x)}{\delta}\right)\right|^{p}\,{\rm d}x\leq C(N,s,p)\frac{r^{2N}}{\lvert\mathcal{A}\rvert}.$
Taking $\delta\rightarrow 0$ in the above estimate, we get $u\equiv 0$ a.e. in $B_{\frac{r}{2}}(x_{j})$. Moreover, $u$ is identically zero on a subset of positive measure in $B_{\frac{r}{2}}(x_{j+1})$ since $\lvert B_{\frac{r}{2}}(x_{j})\cap B_{\frac{r}{2}}(x_{j+1})\rvert>0$. Consequently, repeating the same arguments, we obtain $u\equiv 0$ a.e. in $B_{\frac{r}{2}}(x_{j+1})$, and then $u\equiv 0$ a.e. in $B_{\frac{r}{2}}(x_{i})$ for every $i=1,\dots,k$. Thus $u\equiv 0$ a.e. in $K$. So for every relatively compact set $K$ in ${\mathbb{R}}^{N}$, either $u\equiv 0$ or $u>0$ holds a.e. in $K$. Moreover, there exists a sequence $(K_{n})$ of compact sets such that $\lvert{\mathbb{R}}^{N}\setminus K_{n}\rvert\rightarrow 0$ as $n\rightarrow\infty$. Therefore, either $u\equiv 0$ or $u>0$ also holds a.e. in ${\mathbb{R}}^{N}$. ∎
In the following proposition, we prove that $\tilde{u}$ is positive on ${\mathbb{R}}^{N}$ and $u_{a}$ is non-negative on ${\mathbb{R}}^{N}$ for small enough $a$.
Proposition 3.8.
Let $p\in(\frac{2N}{N+2s},\frac{N}{s})$ and $\tilde{u}$ be given in Proposition 3.5. Then the following hold:
  1. (i)
    $\tilde{u}>0$ a.e. in ${\mathbb{R}}^{N}$.
  2. (ii)
    Let $a_{2}$ be given in Proposition 3.4. Then there exists $a_{3}\in(0,a_{2})$ such that $u_{a}\geq 0$ a.e. in ${\mathbb{R}}^{N}$ for every $a\in(0,a_{3})$.
Proof.
(i) First we show that $\tilde{u}$ is non-negative on ${\mathbb{R}}^{N}$. Consider $A:=\{x\in{\mathbb{R}}^{N}:\tilde{u}\geq 0\}$. Since $\tilde{u}$ is a weak solution of $(-\Delta)_{p}^{s}u=g(x)f_{0}(u)~{}\text{in}~{}\mathcal{D}^{s,p}(\mathbb{R}^{N}),$ we have
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}\tilde{u}(x)-\tilde{u}(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$ $=\int_{{\mathbb{R}}^{N}}g(x)f_{0}(\tilde{u})\phi(x)\,{\rm d}x,$
$=\int_{A}g(x)f_{0}(\tilde{u})\phi(x)\,{\rm d}x,~{}~{}\forall\,\phi\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})},$
where the last identity holds since $f_{0}(t)=0$ for $t\leq 0$. Now we choose $\phi$ to be $-(\tilde{u})^{-}$ to get
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}\tilde{u}(x)-\tilde{u}(y)\big{)}\left((\tilde{u})^{-}(y)-(\tilde{u})^{-}(x)\right)}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=0.$
Further, using ($\bf i{3}$), Remark 3.2, and ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p^{*}_{s}}({\mathbb{R}}^{N})$, we see that
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}\tilde{u}(x)-\tilde{u}(y)\big{)}\left((\tilde{u})^{-}(y)-(\tilde{u})^{-}(x)\right)}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$ $\geq C(p)\iint_{{\mathbb{R}}^{2N}}\frac{\left|(\tilde{u})^{-}(x)-(\tilde{u})^{-}(y)\right|^{p}}{\lvert x-y\rvert^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}$
$\geq C(N,s,p)\left\lVert(\tilde{u})^{-}\right\rVert_{p^{*}_{s}}^{p}.$
Therefore, $(\tilde{u})^{-}=0$ a.e. in ${\mathbb{R}}^{N}$ which implies that $\tilde{u}\geq 0$ a.e. in ${\mathbb{R}}^{N}$. Now we show the positivity of $\tilde{u}$ on ${\mathbb{R}}^{N}$. For a sequence $\{a_{n}\}$ given in Proposition 3.5, since $u_{a_{n}}\rightarrow\tilde{u}$ in $L^{\infty}({\mathbb{R}}^{N})$ (Proposition 3.6) and $\left\lVert u_{a_{n}}\right\rVert_{\infty}\geq C_{2}$ for every large enough $n$ (Proposition 3.4), there exists $C_{3}>0$ such that $\left\lVert\tilde{u}\right\rVert_{\infty}\geq C_{3}$. Now we apply the strong maximum principle (Proposition 3.7) to conclude that $\tilde{u}>0$ a.e. in ${\mathbb{R}}^{N}$.
(ii) Again using $u_{a_{n}}\rightarrow\tilde{u}$ in $L^{\infty}({\mathbb{R}}^{N})$ and $\tilde{u}>0$ a.e. in ${\mathbb{R}}^{N}$, there exists $n_{2}\in{\mathbb{N}}$ such that $u_{a_{n}}\geq 0$ a.e. in ${\mathbb{R}}^{N}$ for all $n\geq n_{2}$. Thus there exists $a_{3}\in(0,a_{2})$ such that for every $a\in(0,a_{3})$, $u_{a}\geq 0$ a.e. in ${\mathbb{R}}^{N}$. ∎
Definition 3.9 (see [11]).
For an open set $\Omega\subset{\mathbb{R}}^{N}$, the space $\widetilde{{\mathcal{D}}}^{s,p}(\Omega)$ is defined as
[…]
It is easy to observe that ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\subset\widetilde{{\mathcal{D}}}^{s,p}(\Omega)$. Let $u\in\widetilde{{\mathcal{D}}}^{s,p}(\Omega)$. We say $(-\Delta)_{p}^{s}u=f$ weakly in $\Omega$, if
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u(x)-u(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}=\int_{\Omega}f(x)\phi(x)\,{\rm d}x,\quad\forall\,\phi\in{\mathcal{C}}_{c}^{\infty}(\Omega).$
First, we recall the following Lemma due to [11, Lemma A.2], which gives an explicit solution on the complement of a ball.
Lemma 3.10.
Let $0<\frac{N-sp}{p-1}<\beta<\frac{N}{p-1}$. For every $R>0$, it holds
$(-\Delta)_{p}^{s}|x|^{-\beta}=C(\beta)|x|^{-\beta(p-1)-sp}\text{ weakly in }\overline{B_{R}}^{c},$
where $C(\beta)$ is given by
$C(\beta)=2\int_{0}^{1}\varrho^{sp-1}\big{[}1-\varrho^{N-sp-\beta(p-1)}\big{]}\big{|}1-\varrho^{\beta}\big{|}^{p-1}\Psi(\varrho)\,\mathrm{d}\varrho,$ (3.36)
and
$\Psi(\varrho)=\mathcal{H}^{N-2}(\mathbb{S}^{N-2})\int_{-1}^{1}\frac{(1-t^{2})^{\frac{N-3}{2}}}{(1-2t\varrho+\varrho^{2})^{\frac{N+sp}{2}}}\,\mathrm{d}t,$ (3.37)
where $\mathcal{H}^{N-2}$ is the Lebesgue measure of dimension $(N-2)$ and $\mathbb{S}^{N-2}$ is a unit sphere in $\mathbb{R}^{N-1}$.
It can be observed from (3.36) that $C(\beta)<0$ since $\beta>\frac{N-sp}{p-1}$. For $u,v\in\widetilde{{\mathcal{D}}}^{s,p}(\Omega)$, we say $(-\Delta)_{p}^{s}u\leq(-\Delta)_{p}^{s}v$ weakly in $\Omega$, if the following holds for all $\phi\in{\mathcal{C}}_{c}^{\infty}(\Omega),\phi\geq 0$:
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u(x)-u(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}\leq\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}v(x)-v(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}.$
Similarly, we say $(-\Delta)_{p}^{s}u\leq(\geq)f$ weakly in $\Omega$, if the following holds for all $\phi\in{\mathcal{C}}_{c}^{\infty}(\Omega),\phi\geq 0$:
$\iint_{{\mathbb{R}}^{2N}}\frac{\Phi\big{(}u(x)-u(y)\big{)}\big{(}\phi(x)-\phi(y)\big{)}}{|x-y|^{N+sp}}\,{\mathrm{d}x\mathrm{d}y}\leq(\geq)\int_{\Omega}f(x)\phi(x)\,{\rm d}x.$
Now we are ready to obtain the positivity of the solution $u_{a}$ for sufficiently small $a$.
Theorem 3.11.
Let $p\in(\frac{2N}{N+2s},\frac{N}{s})$. For $\frac{N-sp}{p-1}<\beta<\frac{N}{p-1}$ assume that $g$ satisfies
$g(x)\leq\frac{B}{|x|^{\beta(p-1)+sp}},\text{ for some constant }B>0\text{ and }x\neq 0.$ (3.38)
Let $u_{a}$ be given in Proposition 3.8. Then there exists $a_{4}\in(0,a_{3})$ such that $u_{a}>0$ a.e. in ${\mathbb{R}}^{N}$ for every $a\in(0,a_{4})$.
Proof.
For $10<\varepsilon<$ 1, we consider the function $\Gamma(z)=\varepsilon^{\beta+1}|z|^{-\beta}$. Then for every $R>0$, $\Gamma\in\widetilde{{\mathcal{D}}}^{s,p}(\overline{B_{R}}^{c})$ and using Lemma 3.10 the following holds weakly:
$(-\Delta)_{p}^{s}\Gamma(z)=\varepsilon^{(\beta+1)(p-1)}C(\beta)|z|^{-\beta(p-1)-sp}\text{ in }\overline{B_{R}}^{c}.$ (3.39)
We define
$\widetilde{\Gamma}(z)=\Gamma(z)-(\Gamma(z)-\varepsilon)_{+}=\min\{\varepsilon,\Gamma(z)\},~{}z\in{\mathbb{R}}^{N}.$
Notice that $\Gamma(z)\geq\varepsilon$ if and only if $|z|\leq\varepsilon$. Thus, the support of the function $(\Gamma(z)-\varepsilon)_{+}$ is contained in the ball $\overline{B_{\varepsilon}}$. Now we choose $x\in\Omega=\overline{B_{R_{1}}}^{c}$ with $R_{1}>2$, $u=\Gamma$, $f=\varepsilon^{(\beta+1)(p-1)}C(\beta)|x|^{-\beta(p-1)-sp}$, $v=-(\Gamma(x)-\epsilon)_{+}$ in [11, Proposition 2.8]. Further, $\widetilde{\Gamma}\in\widetilde{{\mathcal{D}}}^{s,p}(\overline{B_{R_{1}}}^{c})$. Hence, in view of (3.39) with $R=R_{1}$, the following holds weakly in $\overline{B_{R_{1}}}^{c}$:
$(-\Delta)_{p}^{s}\widetilde{\Gamma}(x)$ $=\varepsilon^{(\beta+1)(p-1)}C(\beta)|x|^{-\beta(p-1)-sp}+2\int_{B_{\varepsilon}}\frac{\Phi(\Gamma(x)-\varepsilon)-\Phi(\Gamma(x)-\Gamma(y))}{|x-y|^{N+sp}}\,{\mathrm{d}y}$
$=\varepsilon^{(\beta+1)(p-1)}C(\beta)|x|^{-\beta(p-1)-sp}+2\int_{B_{\varepsilon}}\frac{\Phi(\Gamma(y)-\Gamma(x))-\Phi(\varepsilon-\Gamma(x))}{|x-y|^{N+sp}}\,{\mathrm{d}y}.$ (3.40)
Since $|x|>R_{1}>2$ and $|y|\leq\varepsilon$, it easily follows that $\Gamma(x)<\Gamma(y)$ and $\Gamma(x)<\varepsilon$. Thus, we have the following estimate
$\Phi(\Gamma(y)-\Gamma(x))-\Phi(\epsilon-\Gamma(x))=(\Gamma(y)-\Gamma(x))^{p-1}-(\epsilon-\Gamma(x))^{p-1}\leq(\Gamma(y)-\Gamma(x))^{p-1}\leq(\Gamma(y))^{p-1}.$
Further,
$|x-y|\geq|x|-|y|\geq|x|-\varepsilon\geq|x|-\frac{|x|}{2}=\frac{|x|}{2}.$
Using the above two estimates in (3), the following holds weakly in $\overline{B_{R_{1}}}^{c}$:
$(-\Delta)_{p}^{s}\widetilde{\Gamma}(x)\leq\varepsilon^{(\beta+1)(p-1)}\frac{C(\beta)}{|x|^{\beta(p-1)+sp}}+\frac{2^{N+sp+1}}{|x|^{N+sp}}\int_{B_{\varepsilon}}(\Gamma(y))^{p-1}\,{\mathrm{d}y},$
where we calculate
$\int_{B_{\varepsilon}}(\Gamma(y))^{p-1}\,{\mathrm{d}y}=\varepsilon^{(\beta+1)(p-1)}\int_{B_{\varepsilon}}|y|^{-\beta(p-1)}\,{\mathrm{d}y}$ $=\sigma(\mathbb{S}^{N-1})\varepsilon^{(\beta+1)(p-1)}\int_{0}^{\varepsilon}r^{N-\beta(p-1)-1}\,\mathrm{d}r$
$=\sigma(\mathbb{S}^{N-1})\varepsilon^{(\beta+1)(p-1)}\frac{{\varepsilon}^{N-\beta(p-1)}}{N-\beta(p-1)},$
where the quantity $\sigma(\mathbb{S}^{N-1})$ denotes the $(N-1)$-dimensional measure of the unit sphere in ${\mathbb{R}}^{N}$. Therefore, for any $10<\varepsilon<$ 1, the following holds weakly in $\overline{B_{R_{1}}}^{c}$:
$(-\Delta)_{p}^{s}\widetilde{\Gamma}(x)$ $\leq\left(\frac{C(\beta)}{|x|^{\beta(p-1)+sp}}+\frac{2^{N+sp+1}\sigma(\mathbb{S}^{N-1}){\varepsilon}^{N-\beta(p-1)}}{|x|^{N+sp}(N-\beta(p-1))}\right)\varepsilon^{(\beta+1)(p-1)}$
$\leq\left(C(\beta)+\frac{2^{N+sp+1}\sigma(\mathbb{S}^{N-1}){\varepsilon}^{N-\beta(p-1)}}{(N-\beta(p-1))}\right)\varepsilon^{(\beta+1)(p-1)}|x|^{-\beta(p-1)-sp}$
$:=C_{1}(\beta,\varepsilon)\varepsilon^{(\beta+1)(p-1)}|x|^{-\beta(p-1)-sp}.$
The second last inequality follows from the fact that $|x|>1$ and $\beta(p-1)+sp<N+sp$. Using the fact that $C(\beta)<0$, now we choose $10<\varepsilon<$ 1 small enough so that
${\varepsilon}<\left(\frac{-C(\beta)(N-\beta(p-1))}{2^{N+sp+1}\sigma(\mathbb{S}^{N-1})}\right)^{\frac{1}{N-\beta(p-1)}}.$ (3.41)
Therefore, for $\varepsilon$ as in (3.41) the following holds weakly
$(-\Delta)_{p}^{s}\widetilde{\Gamma}(x)\leq C_{1}(\beta,\varepsilon){\varepsilon}^{(\beta+1)(p-1)}|x|^{-\beta(p-1)-sp}\text{ in }\overline{B_{R_{1}}}^{c},$ (3.42)
where $C_{1}(\beta,\varepsilon)<0$. Since $a_{n}\rightarrow 0$, there exists $n_{1}\in{\mathbb{N}}$ such that $a_{n}B\leq-C_{1}(\beta,\varepsilon){\varepsilon}^{(\beta+1)(p-1)}$ and $u_{a_{n}}\geq 0$ a.e. in ${\mathbb{R}}^{N}$ (Proposition 3.8) for all $n\geq n_{1}$. Thus, for all $n\geq n_{1}$ using the assumption (3.38), the following holds weakly in $\overline{B_{R_{1}}}^{c}$:
$(-\Delta)_{p}^{s}u_{a_{n}}$ $=g(x)(f(u_{a_{n}})-a_{n})\geq-a_{n}g(x)$
$\geq-a_{n}B|x|^{-\beta(p-1)-sp}\geq C_{1}(\beta,\varepsilon){\varepsilon}^{(\beta+1)(p-1)}|x|^{-\beta(p-1)-sp}.$ (3.43)
Further, $u_{a_{n}}\in\widetilde{{\mathcal{D}}}^{s,p}(\overline{B_{R_{1}}}^{c})$. Combining (3.42) and (3), for all $n\geq n_{1}$, the following holds weakly:
$(-\Delta)_{p}^{s}\widetilde{\Gamma}\leq(-\Delta)_{p}^{s}u_{a_{n}}\text{ in }\overline{B_{R_{1}}}^{c}.$ (3.44)
Since $u_{a_{n}}\rightarrow\tilde{u}$ in $L^{\infty}({\mathbb{R}}^{N})$ (Proposition 3.6) and $\tilde{u}>0$ in ${\mathbb{R}}^{N}$ (Proposition 3.8), there exists $\eta_{1}>0$ depending on $R_{1}$ and $n_{2}\in{\mathbb{N}}$ such that $u_{a_{n}}\geq\eta_{1}$ on $\overline{B_{R_{1}}}$ for all $n\geq n_{2}$. For $10<\varepsilon<$ 1 as in (3.41) we further choose $\varepsilon<\eta_{1}$. Hence
$u_{a_{n}}\geq\eta_{1}>\varepsilon\geq\widetilde{\Gamma}\text{ in}\overline{B_{R_{1}}},\quad\forall\,n\geq n_{2}.$ (3.45)
Using (3.44), (3.45), and applying the comparison principle [11, Theorem 2.7], we obtain
$u_{a_{n}}\geq\widetilde{\Gamma}\text{ in }\overline{B_{R_{1}}}^{c},\quad\forall\,n\geq n_{3}:=\max\{n_{1},n_{2}\}.$ (3.46)
Thus, we deduce from (3.45) and (3.46) that $u_{a_{n}}(x)\geq\widetilde{\Gamma}(x)\text{ a.e. in }{\mathbb{R}}^{N}$ for all $n\geq n_{3}$. As a result, there exists $a_{4}\in(0,a_{3})$ such that $u_{a}>0$ a.e. in ${\mathbb{R}}^{N}$ for all $a\in(0,a_{4})$. This completes the proof. ∎
Proof of Theorem 1.1: The proof of part $\rm{(a)}$ follows by Theorem 3.1. Part $\rm{(b)}$ is a consequence of Proposition 3.3. The proof of part $\rm{(c)}$ is provided in Proposition 3.8, whereas the positivity of solutions is demonstrated in Theorem 3.11, which corresponds to part $\rm{(d)}$. ∎
Example 3.12.
Let $N>sp$. For some given constants $A,B>0$ and $\gamma\in(p,p_{s}^{*})$, we consider the following functions:
$f(t)=At^{\gamma-1},\text{ for }t\in{\mathbb{R}}^{+}\text{ and }g(x)=\frac{B}{1+|x|^{\beta(p-1)+sp}},\text{ for }x\in{\mathbb{R}}^{N}.$
$\rm{(i)}$ It is evident that the function $f$ satisfies ($\bf f{1}$), ($\bf f{2}$), and ($\bf\tilde{f{1}}$).
$\rm{(ii)}$ Clearly, $g\in L^{\infty}({\mathbb{R}}^{N})$ and satisfies (1.4). Now we show that $g\in L^{1}({\mathbb{R}}^{N})$.
$\int_{{\mathbb{R}}^{N}}g(x)\,{\rm d}x$ $=\int_{{\mathbb{R}}^{N}}\frac{B}{1+|x|^{\beta(p-1)+sp}}\,{\rm d}x=B\sigma(\mathbb{S}^{N-1})\bigg{(}\int_{0}^{1}+\int_{1}^{\infty}\bigg{)}\frac{t^{N-1}}{1+t^{\beta(p-1)+sp}}\,\mathrm{d}t$
$\leq B\sigma(\mathbb{S}^{N-1})\bigg{(}1+\int_{1}^{\infty}t^{N-1-\beta(p-1)-sp}\,\mathrm{d}t\bigg{)}\leq C(N,s,p,\beta,B).$
The last integral is finite since $\beta(p-1)>N-sp$.
Acknowledgements: N. Biswas is funded by the Department of Atomic Energy, Government of India, under project no. 12-R&\&&D-TFR-5.01-0520 and the Science and Engineering Research Board, Government of India, through the National Postdoctoral Fellowship (PDF/2023/000038). R. Kumar acknowledges the support of the CSIR fellowship, file no. 09/1125(0016)/2020–EMR–I. The authors thank Prof. Abhishek Sarkar (IIT Jodhpur) for the beneficial discussions.
Declarations: The authors declare that there is no conflict of interest. Data sharing does not apply to this article, as no data sets were generated or analyzed during the current study.
Appendix A
This section contains some technical results.
Lemma A.1.
Let $p\in(1,\infty),~{}\beta\geq 1,a,b\in{\mathbb{R}}$ and $M>0$. Let $C(\beta,p)=\big{(}\frac{p}{\beta+p-1}\big{)}^{p}$. Then the following inequalities hold:
  1. ($\bf i{1}$)
    $|a-b|^{p-2}(a-b)\left((a^{+})^{\beta}-(b^{+})^{\beta}\right)\geq C(\beta,p)\left|(a^{+})^{\frac{\beta+p-1}{p}}-(b^{+})^{\frac{\beta+p-1}{p}}\right|^{p}$ where $a^{+}=\max\{a,0\}$.
  2. ($\bf i{2}$)
    $|a-b|^{p-2}(a-b)\left((a^{+})_{M}^{\beta}-(b^{+})_{M}^{\beta}\right)\geq C(\beta,p)\left|(a^{+})_{M}^{\frac{\beta+p-1}{p}}-(b^{+})_{M}^{\frac{\beta+p-1}{p}}\right|^{p}$ where $(a^{+})_{M}=\min\{a^{+},M\}$ and $(b^{+})_{M}=\min\{b^{+},M\}$.
  3. ($\bf i{3}$)
    $|a-b|^{p-2}(a-b)\left((b^{-})^{\beta}-(a^{-})^{\beta}\right)\geq C(\beta,p)\left|(a^{-})^{\frac{\beta+p-1}{p}}-(b^{-})^{\frac{\beta+p-1}{p}}\right|^{p}$ where $a^{-}=-\min\{a,0\}$.
  4. ($\bf i{4}$)
    $|a-b|^{p-2}(a-b)\left((b^{-})_{M}^{\beta}-(a^{-})_{M}^{\beta}\right)\geq C(\beta,p)\left|(a^{-})_{M}^{\frac{\beta+p-1}{p}}-(b^{-})_{M}^{\frac{\beta+p-1}{p}}\right|^{p}$ where $(a^{-})_{M}=\min\{a^{-},M\}$ and $(b^{-})_{M}=\min\{b^{-},M\}$.
Proof.
(i) If $a=b$, then ($\bf i{1}$) holds trivially. So we assume that $a\neq b$. Without loss of generality, we can assume that $a>b$. For $a,b>0$, $a^{+}=a,b^{+}=b$, and ($\bf i{1}$) follows using [9, Lemma C.1]. If $a,b<0$ then ($\bf i{1}$) holds trivially since $a^{+}=0=b^{+}$. Next, we assume $a>0>b$. Since $a^{+}=a$ and $b^{+}=0$, we need to show that
$(a-b)^{p-1}\geq C(\beta,p)a^{p-1}.$ (A.1)
If we divide (A.1) by $a^{p-1}$, we get $(1-\frac{b}{a})^{p-1}\geq C(\beta,p)$ and this inequality always holds true since $C(\beta,p)\leq 1$ and $\frac{b}{a}<0$.
(ii) Now we consider $M>0$. For $a=b$, ($\bf i{2}$) holds trivially. So without loss of generality, we assume that $a>b$. If $a,b\geq M$ then $(a^{+})_{M}=M=(b^{+})_{M}$, and ($\bf i{2}$) holds trivially. If $a,b\leq M$, then by noting that $(a^{+})_{M}=a^{+},(b^{+})_{M}=b^{+}$, ($\bf i{2}$) follows using ($\bf i{1}$). Now we assume that $b<M<a$. In this case $(b^{+})_{M}=b^{+}<M=(a^{+})_{M}$. Hence using ($\bf i{1}$) we get
$|a-b|^{p-2}(a-b)\big{(}(a^{+})_{M}^{\beta}-(b^{+})_{M}^{\beta}\big{)}$ $\geq|M-b|^{p-2}(M-b)\left(M^{\beta}-(b^{+})^{\beta}\right)$
$\geq C(\beta,p)\left|M^{\frac{\beta+p-1}{p}}-(b^{+})^{\frac{\beta+p-1}{p}}\right|^{p}$
$=C(\beta,p)\left|(a^{+})_{M}^{\frac{\beta+p-1}{p}}-(b^{+})_{M}^{\frac{\beta+p-1}{p}}\right|^{p}.$
Thus ($\bf i{2}$) holds for every $a,b\in{\mathbb{R}}$.
(iii) Without loss of generality, assume that $a>b$. If $a,b>0$ then ($\bf i{3}$) holds trivially since $a^{-}=0=b^{-}$. For $a,b<0$, $a^{-}=-a,b^{-}=-b$, where $0\leq a^{-}<b^{-}$. Applying [9, Lemma C.1] we get
$|a^{-}-b^{-}|^{p-2}(a^{-}-b^{-})\left((a^{-})^{\beta}-(b^{-})^{\beta}\right)\geq C(\beta,p)\left|(a^{-})^{\frac{\beta+p-1}{p}}-(b^{-})^{\frac{\beta+p-1}{p}}\right|^{p}.$
The above inequality infers ($\bf i{3}$). Next, we consider $a>0>b$. In this case, ($\bf i{3}$) has the following form:
$(a-b)^{p-1}\geq C(\beta,p)(-b)^{p-1}.$
Dividing the above inequality by $(-b)^{p-1}$, we see $\left(-\frac{a}{b}+1\right)^{p-1}\geq C(\beta,p)$. This inequality always holds true since $-\frac{a}{b}\geq 0$ and $C(\beta,p)\leq 1$.
(iv) Now we consider $M>0$. For $a=b$, ($\bf i{4}$) holds trivially. So without loss of generality, we assume that $a>b$. For $a,b\geq 0$, ($\bf i{4}$) trivially holds. If $a,b\leq-M$, then $(a^{-})_{M}=M=(b^{-})_{M}$, and ($\bf i{4}$) holds. If $a,b\geq-M$, then by noticing that $(a^{-})_{M}=a^{-},(b^{-})_{M}=b^{-}$, ($\bf i{4}$) follows using ($\bf i{3}$). Now we assume that $b<-M<a$. For the case $a,b<0$, we notice that $(a^{-})_{M}=a^{-},(b^{-})_{M}=M$ and consequently, using ($\bf i{3}$) we obtain
$|a-b|^{p-2}(a-b)\big{(}(b^{-})_{M}^{\beta}-(a^{-})_{M}^{\beta}\big{)}$ $\geq|a-(-M)|^{p-2}(a-(-M))\big{(}(M)^{\beta}-(a^{-})^{\beta}\big{)}$
$\geq C(\beta,p)\bigg{|}(a^{-})^{\frac{\beta+p-1}{p}}-M^{\frac{\beta+p-1}{p}}\bigg{|}^{p}$
$=C(\beta,p)\bigg{|}(a^{-})_{M}^{\frac{\beta+p-1}{p}}-(b^{-})_{M}^{\frac{\beta+p-1}{p}}\bigg{|}^{p}.$
Now if we consider $a>0>b$, then $(a^{-})_{M}=0,(b^{-})_{M}=M$. In this case,
(i4)$\Longleftrightarrow(a-b)^{p-1}(b^{-})_{M}^{\beta}\geq C(\beta,p)(b^{-})_{M}^{\beta+p-1}\Longleftrightarrow\bigg{(}\frac{a-b}{M}\bigg{)}^{p-1}\geq C(\beta,p).$
The last inequality holds since $C(\beta,p)\leq 1$ and $\frac{a}{M}-\frac{b}{M}>\frac{a}{M}+1>1$. Thus, ($\bf i{4}$) holds for every $a,b\in{\mathbb{R}}$. ∎
Lemma A.2.
An iterative sequence is defined as
$l_{0}=p_{s}^{*},\quad l_{j+1}=\theta l_{j}+\frac{p^{*}_{s}(p-1)}{p},\text{ and }\theta=\frac{p^{*}_{s}}{p\bar{r}^{\prime}},$ (A.2)
where $\bar{r}^{\prime}=\frac{\bar{r}}{\bar{r}-1}$ and $\bar{r}$ is given by (3.23). Then there exist $\beta_{1},\beta_{2}>0$ such that $\beta_{1}<\frac{\theta^{j}}{l_{j}}<\beta_{2}$ for all $j\in{\mathbb{N}}$.
Proof.
From (A.2), we can write the iterative sequence as
$l_{j}=\theta^{j}l_{0}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=0}^{j-1}\theta^{i}=\theta^{j}p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=0}^{j-1}\theta^{i}.$ (A.3)
Dividing (A.3) by $\theta^{j}$, we obtain
$\frac{l_{j}}{\theta^{j}}$ $=p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\left(\frac{1}{\theta^{j}}+\frac{1}{\theta^{j-1}}+\cdots+\frac{1}{\theta}\right)=p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=1}^{j}\frac{1}{\theta^{i}}<p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=1}^{\infty}\frac{1}{\theta^{i}}.$
Consequently,
$\frac{l_{j}}{\theta^{j}}<p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\frac{1}{\theta-1}.$ (A.4)
Furthermore, it is evident from (A.3) that
$\frac{l_{j}}{\theta^{j}}>p_{s}^{*}.$ (A.5)
The combination of (A.4) and (A.5) yields $\beta_{1}<\frac{\theta^{j}}{l_{j}}<\beta_{2},$ for all $j\in{\mathbb{N}}$, where $\beta_{1}=\frac{1}{p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p(\theta-1)}}$ and $\beta_{2}=\frac{1}{p_{s}^{*}}$. ∎